### v1

# GroupNL: Low-Resource and Robust CNN Design over Cloud and Device

Abstract
It has become mainstream to deploy Convolutional Neural Network (CNN) models on ubiquitous Internet of Things (IoT) devices with the help of the cloud to provide users with a variety of high-quality services. Most existing methods have two limitations: (i) low robustness in handling corrupted image data collected by IoT devices; and (ii) high consumption of computational and transmission resources. To this end, we propose the Grouped NonLinear transformation generation method (GroupNL), which generates diversified feature maps by utilizing data-agnostic Nonlinear Transformation Functions (NLFs) to improve the robustness of the CNN model. Specifically, partial convolution filters are designated as seed filters in a convolutional layer, and a small set of feature maps, i.e., seed feature maps, are first generated based on vanilla convolution operation. Then, we split seed feature maps into several groups, each with a set of different NLFs, to generate corresponding diverse feature maps with in-place nonlinear processing. Moreover, GroupNL effectively reduces the parameter transmission between multiple nodes during model training by setting the hyperparameters of NLFs to random initialization and not updating them during model training, and reduces the computing resources by using NLFs to generate feature maps instead of most feature maps generated based on sliding windows. Experimental results on CIFAR-10, GTSRB, CIFAR-10-C, Icons50, and ImageNet-1K datasets in NVIDIA RTX GPU platforms show that the proposed GroupNL outperforms other state-of-the-art methods in model robust and training acceleration. Specifically, on the Icons-50 dataset, the accuracy of GroupNL-ResNet-18 achieves approximately 2.86% higher than the vanilla ResNet-18. GroupNL improves training speed by about 53% compared to vanilla CNN when trained on a cluster of 8 NVIDIA RTX 4090 GPUs on the ImageNet-1K dataset.
Index Terms:
Internet of Things, cloud computing, cloud-assisted, CNNs.
1 Introduction
Internet of Things (IoT) Analytics predicts the number of connected IoT devices worldwide will increase from 16.7 billion in 2023 to 29 billion in 2027111https://iot-analytics.com/number-connected-iot-devices/. It is possible to provide convenient services with the help of connected and ubiquitous IoT devices. Additionally, given the excellent performance of convolutional neural networks (CNN) [1, 2, 3, 4] in computer vision applications, it has become mainstream to deploy CNN models on IoT devices to provide people with various convenient and high-quality services, such as image recognition and monitoring services [5, 6, 7, 8, 9].
Figure 1: Overview of cloud-assisted architecture.
IoT devices such as smart cameras and monitors usually work outdoors and often need to process data affected by the natural environment (such as fog, rain, snow, etc.). Vanilla CNN models have low robustness when processing such environmentally affected data. In addition, IoT devices have limited computing and storage resources. Using cloud servers to train CNN models and then sending them to IoT devices for deployment and service provision has become mainstream, as shown in Fig. 1. However, a cloud server usually corresponds to thousands of IoT devices, and frequent deployment and updating of vanilla CNN models will occupy a lot of bandwidth resources. In addition, the vanilla CNN generates feature maps through sliding windows, which consumes a lot of floating-point operations (FLOPs). Therefore, to provide high-quality services on IoT devices, it is urgent to design a CNN model that is highly robust, parameter transmission-friendly and has low computing resource requirements.
This paper aims to propose a CNN design method that increases the diversity of feature maps to improve model robustness without increasing model complexity while ensuring parameter transmission friendliness when deploying and updating models for cloud-assisted IoT devices.
To this end, first, this paper proposes the Grouped NonLinear transformation method, GroupNL. In the GroupNL-based CNN model, partial trainable filters are configured within each convolution layer, which is referred to as seed filters. The feature maps produced via seed filters are called seed feature maps. Subsequently, we split the seed feature maps into many small groups. The remaining feature maps can be generated with diverse nonlinear transformation functions (NLFs) configured with different hyperparameters for each group of seed feature maps. The seed feature maps and generated feature maps are concatenated as the output of the convolutional layer.
GroupNL has four advantages: (i) By grouping seed feature maps, GroupNL can generate diversified feature maps using NLFs with different hyperparameters without increasing the complexity of the model, effectively improving the robustness of the model. (ii) The hyperparameters of NLFs are randomly initialized and remain unchanged during model training. They can be saved and reproduced using random numbers so that after the cloud trains the model, only a small amount of learning parameters and random numbers need to be transmitted. Compared with transmitting the complete model, the amount of parameters transmitted to the IoT device can be effectively reduced. In addition, when the GroupNL-based CNN model employing parallel training technologies [10, 11, 12] like data parallelism (DP) and distributed data parallelism (DDP) to accelerate model training, only a small number of gradients in seed filters have to exchange between multi-GPUs for scatter and gather. This process reduces the intermediate data exchange between different GPUs. Compared with scattering and gathering all gradients in the vanilla CNN model, only exchanging a small number of gradients from seed filters in multi-GPUs memory dramatically accelerates the training speed. (iii) GroupNL includes generating feature maps based on sliding windows and nonlinear transformations. Compared with the vanilla CNN that generates feature maps completely based on sliding windows, it can effectively reduce the number of FLOPs required to run the CNN model. (iv) GroupNL is simple and easy to use, and the convolution modules it generates can directly replace the convolution layers in the existing CNN architecture.
Then, for larger CNN models, such as ResNet-101, this paper introduces sparse technology, that is, apply group convolution to generate seed feature maps, which further reduces the number of learnable parameters and FLOPs of the GroupNL-based model while ensuring model performance. This paper also analyzes four types of NLFs, i.e., Sinusoidal, Monomial, Gaussian, and Laplace functions. It verifies on multiple datasets that the feature maps generated based on the Sinusoidal function have the best performance.
Finally, this paper conducts extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, Icons-50, and ImageNet-1K datasets, and experimental results demonstrate that our proposed GroupNL-based CNN models achieve strong performance in robustness and resource-saving. For example, the accuracy of GroupNL ResNet-34 on CIFAR-10-C is about 2.0% higher than that of vanilla ResNet-34, and the accuracy of GroupNL ResNet-18 on Icons-50 is about 2.86% higher than that of vanilla ResNet-18. We also verify the training acceleration in CIFAR-10 and ImageNet-1K datasets on the NVIDIA RTX 2080Ti/RTX 4090 multi-GPUs platforms. Particularly, with the distributed data parallel training in ImageNet-1K, the GroupNL ResNet-101 achieves an appropriately 53% speedup compared to the vanilla ResNet-101 while maintaining accuracy.
In summary, our main contributions are as follows.
  • This paper proposes the GroupNL, which groups seed feature maps and generates the remaining feature maps using the NLF with different hyperparameters for each group. This improves the diversity of generated feature maps and the robustness of the CNN model without increasing its complexity.
  • This paper introduces sparse technology to the ResNet-101 model, further reducing the number of learnable parameters and FLOPs of the model while maintaining model performance. In addition, it verifies that using the Sinusoidal function is the best to generate feature maps for GroupNL.
  • Extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, and Icons-50, and ImageNet-1K datasets demonstrate that our proposed GroupNL achieves strong performance in terms of model robustness, number of learnable parameters, FLOPs, and training speedup.
The remainder of the paper is organized as follows. Section 2 reviews the work on CNN and cloud-assisted CNN training and deployment. Section 3 describes the proposed GroupNL method in detail. Section 4 presents our evaluation results, and Section 5 concludes the paper.
2 Related Work
This section briefly introduces the Convolutional Neural Network (CNN) models and cloud-device CNN models training and deployment, which are inspiring our work.
2.1 Convolutional Neural Network Models
Because of the excellent feature extraction capabilities of the convolution operation, many popular CNN models based on convolution operations have been proposed. For example, Simonyan et al. [1] proposed the VGG model through stacked convolution operations with multiple model versions in different layers. He et al. [2] proposed the ResNet with residual branching, which solves the problem of degradation caused by deepening the convolutional layers. Ding et al. [4] introduced the re-parameterization technology to speed up the inference of the CNN model. To deploy CNN models in resource-constrained IoT devices, many lightweight CNN models have been proposed, such as MobileNet [13, 14, 15], ShuffleNet [16, 17], and GhostNet [18], etc. Some methods also consider the robustness of the CNN model, such as MonoCNN [5] and SineFM [19] for cloud-device collaborative deployment.
TABLE I: Feature Maps Generated-based CNN Methods Training Evaluation with ResNet-101 in CIFAR-10 for 300 Epochs on NVIDIA RTX 2080Ti GPUs.
Method Full Conv Nonlinear Func Cheap Linear Norm Top-1 Acc (%) Training Time
Vanilla CNN 95.48 9.3 Hours
MonoCNN 94.97 9.2 Hours
GhostNet 95.04 12.7 Hours
SineFM 95.06 15.4 Hours
GroupNL (Ours) 95.51 8.3 Hours
The most similar to our proposed GroupNL method are the GhostNet [18] and SineFM [19], which also generate feature maps via transformation. However, there are two significant differences as shown in Table I:
(i) The GhostNet generates the feature maps with cheap linear transformation. The cheap linear is a variant of depthwise convolution (DWConv), in which the number of output channels is the multiple of the number of input channels, and the groups of convolution are equal to the number of input channels. So in this paper, we still refer to the cheap linear transformation as “DWConv”. Introducing trainable nodes, i.e., DWConv in the computational graph will result in slower training than vanilla CNN models.
(ii) The SineFM mixes the nonlinear and cheap linear transformation for the feature maps generation, and also introduces batch normalization (Norm) for generated feature maps; SineFM has a better performance compared to GhostNet via nonlinear, nevertheless, it introduces trainable Norm and DWConv resulting in slower training.
Compared to the above methods, our proposed GroupNL is designed with grouped nonlinear transformation to ensure the diversity of generated feature maps and better performance. At the same time, the generation in GroupNL only contains the copying and in-placing operators without extra trainable computational nodes, e.g., DWConv and Norm.
2.2 Cloud-Device CNN Training and Deployment
Due to the limited resources of IoT devices, cloud-device collaborative training and deployment of CNN models have received widespread attention. The existing architectures can be classified into three categories:
Cloud-Only. The cloud-only architecture is training and deploying the CNN models on the cloud server. The user sends data and service requests to the cloud server, and after the cloud server processes the data, it sends the results to the users [20, 21, 22, 23]. For example, Liu et al. [23] proposed to deploy the CNN model on the cloud server, and the device obtains recognition services by uploading the collected data to the cloud server. They also process data on edge servers [24, 25, 26, 27, 28, 29] to reduce the amount of data uploaded to the cloud. However, the quality of services of cloud-only architecture is fully constrained by the network condition.
Cloud-Device Collaboration. The second architecture splits the CNN model into two parts, partial CNN layers are deployed on the cloud server and other parts on the IoT devices. The cloud and IoT devices are collaborative training and inference of the split CNN model [30, 31, 32]. For example, Kang et al. [30] proposed to divide the CNN into a head running on the device and a tail running on the cloud and decide the split point based on the load of the device and the conditions of the cloud and model. Stefanos et al. [32] proposed a cloud-device collaborative inference method for the cloud and device with compression technologies to reduce the parameters exchange. However, cloud-device collaborative CNN methods have to search the optimal partitioning points labor-intensively for different CNN models.
Cloud-Assisted Training and Device Deployment. The third architecture is that the cloud server first trains the CNN model and then sends the trained CNN model to the IoT device for deployment [5, 6, 7, 33, 34]. We refer to it as the Cloud-Assisted Architecture. For example, Ding et al. [5] proposed to train a CNN model containing a few learnable parameters on the cloud server and then send the trained CNN model to the IoT device for deployment. Lu et al. [34] proposed to train a lightweight vision transformer model on the cloud server, and then send the trained vision transformer model to the IoT devices for deployment. Furthermore, they reduce the number of model parameters transmitted from the cloud to IoT devices by replacing the attention mechanism with multi-scale pooling technology.
The first and second architectures are closely related to network conditions. When the network connection is unstable or unavailable, IoT devices will provide poor-quality services or even be unable to provide services. The proposed GroupNL is based on the cloud-assisted training and IoT device deployment architecture.
3 Design of the Proposed Approach
3.1 Overview
The proposed GroupNL is based on the cloud-assisted IoT device deployment and updating CNN model architecture. In this architecture, the CNN model is first trained in the cloud server, and then the trained CNN model is transferred to the IoT device for deployment and visual service.
Two key factors motivate us to investigate a robust and low-resource CNN design method.
(i) Typically, IoT devices such as smart surveillance and autonomous driving cameras work outdoors, and the CNN models deployed on the devices often process corrupted data affected by the natural environment, such as images collected on rainy, foggy, and snowy days. We refer to the performance of the on-device CNN model on corrupted data as model robustness. Most existing CNN models have low robustness when dealing with corrupted data.
(ii) IoT devices are resource-constrained. To address the high resource consumption of vanilla CNN models, many alternatives to CNN have been proposed. Although alternatives to the vanilla CNN model reduce the number of model parameters and computation, they contain many additional operations (e.g., DWConv, Norm, etc.), which results in no significant acceleration of model training speed.
To this end, we aim to design a novel, robust, and low-resource CNN model design method to improve the quality of service provided to users by cloud-assisted IoT devices in deploying and updating CNNs.
Figure 2: An example of the GroupNL operation with $c_{\mathrm{in}}=8,c_{\mathrm{out}}=16$, and the number of seed features is $c_{\mathrm{seed}}=4$, the groups for splitting is $g=2$. In this case, the GroupNL can introduce $\gamma\cdot g=6$ different NLFs for the feature maps generation. Compared with the vanilla convolution operation needs 16 filters to compute all features, the GroupNL operation only requires 4 filters to compute the seed features and generates the remaining features by splitting, repeating, and diverse nonlinear transformations.
3.2 Preliminaries of Nonlinear Transformation and Feature Maps Generated-based Methods in CNNs
We illustrate our motivation by analyzing the most relevant works, nonlinear transformation method, i.e., MonoCNN [5], and feature maps generated-based methods, namely GhostNet [18] and SineFM [19], respectively. We also describe the operations of these CNN design methods.
Given the input feature maps ${\bm{x}}$ and output feature maps ${\bm{y}}$ of this convolutional layer, where ${\bm{x}}\in\mathbb{R}^{h\times w\times c_{\mathrm{in}}}$ and ${\bm{y}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times w^{\!{}^{\prime}\!}\times c_{\mathrm{out}}}$. The $h$ and $w$ are the spatial dimensions of the input data/feature maps and $h^{{}^{\prime}}$ and $w^{{}^{\prime}}$ are the spatial dimensions of the output feature maps. $c_{\mathrm{in}}$ is the number of input channels and $c_{\mathrm{out}}$ is the number of output channels.
For a detailed description of the nonlinear transformation and feature maps generated-based methods, we first introduce two concepts from these CNNs:
(i) seed filters refer to a small number of convolutional filters in the convolution layer.
(ii) seed feature maps refer to the feature maps generated by the seed filters and input via a convolution operation.
We denote the seed filters are the ${\bm{W}}_{\mathrm{seed}}$, where ${\bm{W}}_{\mathrm{seed}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{seed}}}$ where $k$ is the spatial dimension of the filter assumed to be square. The number of seed filters is $c_{\mathrm{seed}}$, and the number of generated filters is $c_{\mathrm{gen}}$, where $c_{\mathrm{seed}}$+$c_{\mathrm{gen}}$ = $c_{\mathrm{out}}$. We set the trainable computational operations belonging to $\tt{torch.nn.Module}$ in bold in equation, i.e., $\mathrm{\bf{DWConv}}$ and $\mathrm{\bf{Norm}}$. We also denote a reduction ratio $r$ representing the relationship between $c_{\mathrm{seed}}$ and $c_{\mathrm{out}}$:
$\begin{array}[]{l}r={c_{\mathrm{out}}}/{c_{\mathrm{seed}}}\end{array}.$ (1)
MonoCNN [5]. For a convolutional layer, some filters are designated as seed filters, and the remaining filters are generated based on the seed filters and the data-agnostic nonlinear transformation function (NLF). Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$, and a monomial NLF $\Psi_{\mathrm{mono}}(\cdot)$, we can obtain the remaining filters ${\bm{W}}_{\mathrm{gen}}$ of this layer as follows:
$\begin{array}[]{l}{\bm{W}}_{\mathrm{gen}}=\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})\end{array},$ (2)
where ${\bm{W}}_{\mathrm{gen}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{gen}}}$. Then, concatenating the seed filter ${\bm{W}}_{\mathrm{seed}}$ and the generated filters ${\bm{W}}_{\mathrm{gen}}$ to obtain the filters ${\bm{W}}$ of this layer:
$\begin{array}[]{l}{\bm{W}}=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};{\bm{W}}_{\mathrm{gen}}]\end{array},$ (3)
where ${\bm{W}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{out}}}$. Finally, perform the vanilla convolution operation on the input data/feature maps to obtain the output ${\bm{y}}$ of this layer:
${\bm{y}}$ $={\bm{W}}*{\bm{x}}$ (4)
$=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})]*{\bm{x}}\ ,$
where * is the convolution operation.
The MonoCNN improves the robustness of the CNN model via data-agnostic NLFs for filters but still with a complete convolution operation.
GhostNet [18]. The GhostNet introduces partial convolution filters ${\bm{W}}_{\mathrm{seed}}$ and generated the seed feature maps as follows:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}={\bm{W}}_{\mathrm{seed}}*{\bm{x}}\end{array},$ (5)
where ${\bm{y}}_{\mathrm{seed}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times c_{\mathrm{seed}}}$. For the remaining feature maps, GhostNet introduces additional cheap linear transformation, i.e., the $d\times d$ depthwise convolution filters to process the ${\bm{y}}_{\mathrm{seed}}$ for generating features:
${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (6)
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{d\times d}({\bm{y}}_{\mathrm{seed}})]\ .$
However, the cheap linear operation (i.e., DWConv op.) introduced by GhostNet causes a slow computation in training. At the same time, the parameters of DWConv are also trainable, so the robustness of GhostNet is completely restricted to the training data.
SineFM [19]. Similar to the GhostNet, some filters of SineFM are designated as seed filters for a convolution layer. First, seed feature maps equal to the number of seed filters are generated based on the seed filters and input data. Then, feature maps with $t$-times the number of $c_{\mathrm{seed}}$ are generated based on the seed feature maps and the NLFs. Finally, using $1\times 1$ depthwise convolution for channel alignment.
Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer by the Eq. 5. Then, $t$-data-agnostic NLFs $\bm{{\Psi}}_{\mathrm{sinefm}}={\{\Psi_{i}\|i=1,\dots,t\}}$ process the seed feature maps to obtain more features:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{exp}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{exp}_{1}};\cdots;{\bm{y}}_{\mathrm{exp}_{t}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[{\bm{y}}_{\mathrm{exp}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{exp}_{i}}=\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))\ ,\end{array}$ (7)
where ${\bm{y}}_{\mathrm{exp}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!c_{\mathrm{seed}}}$, ${\bm{y}}_{\mathrm{exp}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\times\!t\cdot c_{\mathrm{seed}}}$, and $i\!=\!1,\!\cdots,\!t$. The $t$ is the times of expanding seed feature maps. The generated feature maps expanding by $t$ times of seed feature maps will usually exceed the number of feature maps required for this layer. Each expanding seed feature maps ${\bm{y}}_{\mathrm{exp}_{i}}$ also process a specific $\mathrm{\bf{Norm}}_{i}$. All expanding seed feature maps are concatenated and then process the $1\times 1$ depthwise convolution filters to align the dimensions and obtain the generated feature maps:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}({\bm{y}}_{\mathrm{exp}})\ .\end{array}$ (8)
Finally, the generated feature maps are concatenated with the seed feature maps to form the output of the SineFM layer as follows:
${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (9)
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}(\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))])]\ .$
3.3 Design of GroupNL
3.3.1 Grouped Nonlinear Transformation Algorithm
The aforementioned GhostNet and SineFM methods introduce too much computation in feature maps generation by DWConv and Norm operations which result in more computational graph nodes and slow down training. At the same time, the data-agnostic nonlinear transformation of SineFM is only $t$-different types.
Therefore, resource efficiency and robustness are needed to be further improved. The key challenge is how to increase the diversity of data-agnostic nonlinear transformations to improve the robustness of the model without increasing the resource computation of running the model.
To this end, we propose the Grouped NonLinear transformation feature map generation (GroupNL) method, which is a robust and low-resource CNN model design method. Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input data/feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer as same as the Eq. 5.
Besides, we designate the seed filters ${\bm{W}}_{\mathrm{seed}}$ as the standard convolution in basic structures like ResNet-18/-34 and VGG11, and sparse in bottleneck structures, e.g., ResNet-101. The sparse filter means that the number of groups of the convolution operation of seed filters is the Greatest Common Divisor (GCD) of $c_{\mathrm{in}}$ and $c_{\mathrm{seed}}$, i.e., $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. Introducing the sparse seed convolution in the larger bottleneck structure can further improve the resource efficiency of the larger CNN models. Utilizing the standard seed convolution in basic structure is beneficial for maintaining the performance of the smaller CNN models.
Splitting and Grouped Algorithm. We split the seed feature maps into $g$ groups in $c_{\mathrm{seed}}$ dimension instead of expanding, as follows:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}_{1}};\cdots;{\bm{y}}_{\mathrm{seed}_{g}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[{\bm{y}}_{\mathrm{seed}_{i}}]\end{array},$ (10)
where $g$ is a positive integer, and the modulo result is $\mathrm{Mod}(c_{\mathrm{seed}},g)\!=\!0$, and ${\bm{y}}_{\mathrm{seed}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\frac{c_{\mathrm{seed}}}{g}},i=1,\cdots,g$. For the remaining dimension $c_{\mathrm{gen}}=c_{\mathrm{out}}-c_{\mathrm{seed}}$, we want to represent it by $c_{\mathrm{seed}}$ without expanding the dimension.
We denote a multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=(r\cdot c_{\mathrm{out}}/c_{\mathrm{in}}-1)$, so each group of seed feature maps should be copied $\gamma$ times for filling the remaining dimension. Thus, the total number of copied groups of feature maps is $\gamma\cdot g$.
Grouped Diverse Nonlinear Transformation. By splitting and copying, we have $\gamma\cdot g$ groups of identical seed feature maps, and no additional dimensions are introduced. We can introduce $\gamma\cdot g$ data-agnostic different nonlinear transformations, i.e., diverse NLFs $\bm{{\Psi}}_{\mathrm{groupnl}}={\{\Psi_{i,j}\|i=1,\dots,g;j=1,\dots,\gamma\}}$ to these identical groups. Subsequently, for each copied group of seed feature maps, NLF with different hyperparameter settings is used to generate the corresponding feature maps, as follows:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{gen}_{1}};\cdots;{\bm{y}}_{\mathrm{gen}_{g}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[{\bm{y}}_{\mathrm{gen}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{gen}_{i}}=\mathop{\mathrm{Concat}}\limits_{j=1,\cdots,\gamma}[\Psi_{i,j}({\bm{y}}_{\mathrm{seed}_{i}})]\ ,\end{array}$ (11)
where $i\!=\!1,\!\cdots,\!g$, $j\!=\!1,\!\cdots,\!\gamma$, ${\bm{y}}_{\mathrm{gen}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\gamma\cdot\frac{c_{\mathrm{seed}}}{g}}$, and ${\bm{y}}_{\mathrm{gen}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\gamma\cdot c_{\mathrm{seed}}},\gamma\cdot c_{\mathrm{seed}}=c_{\mathrm{gen}}$.
As shown in Fig. 2, we can describe the operations of grouped diverse nonlinear transformation in detail with an example of $c_{\mathrm{in}}=8,c_{\mathrm{out}}=16,c_{\mathrm{seed}}=4,g=2$. In this case, the number of generated features $c_{\mathrm{gen}}=16-4=12$, the multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=12/4=3$, and introducing $|\bm{{\Psi}}_{\mathrm{groupnl}}|=\gamma\cdot g=6$ different NLFs for feature maps generation.
Finally, generated feature maps are concatenated with the seed feature maps to produce the output of the GroupNL layer, as follows:
${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (12)
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[\mathop{\mathrm{Concat}}\limits_{j=1,\cdots,\gamma}[\Psi_{i,j}({\bm{y}}_{\mathrm{seed}_{i}})]]]\ .$
The PyTorch-like pseudocode of the GroupNL-based convolutional layer is demonstrated in Algorithm 1.
As described in Eq. 12 and Algorithm 1, the generating procedure of GroupNL is all in-placing (e.g., splitting $\tt{torch.split}$, concatenation $\tt{torch.cat}$) and copying (e.g., $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$) operators, and data-agnostic and random nonlinear transformation operations, without any extra $\tt{torch.nn.Module}$ (i.e., DWConv and Norm) compared with the GhostNet in Eq. 6 and SineFM in Eq. 9.
# func: nonlinear transformation function
# sparse: is seed filter sparse
# g: No. of nonlinear transformation groups
# C_i, C_s, C_o: No. of input/seed/output features
class GroupNLConv2d(nn.Conv2d):
def __init__(self, C_in, C_out, C_s, g, func, sparse, **kwargs):
super(GroupNLConv2d, self).__init__(
C_in, C_out, **kwargs)
self.weight = None # ensure non-learnable
self.sparse = sparse # sparsity flag
# No. of generated features
C_g = C_s * (ceil(C_o/C_s) - 1)
# seed filters for generating feasures
self.conv = nn.Conv2d(C_i, C_s, \
groups=math.gcd(C_i, C_s) \
if self.sparse else 1, **kwargs)
# No. of grouped seed features lists
groups_s = [C_s//g for _ in range(g)]
gamma = C_g/C_s # multiplier gamma
# diverse random hyperparameters of func
self.hy=nn.Parameter(torch.rand(gamma*g))
# ensure non-learnable hyperparameters
self.hy.requires_grad = False
def forward(self, x):
# feature maps shape [B, C, H, W]
y_s = self.conv(x) # seed features
# split seed features into groups
y_sp = torch.split(y_s, groups_s)
# copy gamma times feature maps
y_cp = torch.cat([y_sp[j].repeat(
[1, gamma, 1, 1]) for j in range(g)])
# copy hyperparameters of func
hy = self.hy.repeat_interleave(C_s/g).reshape([1, C_g, 1, 1])
y_g = func(y_cp, hy) # generate C_g features
return torch.cat([y_s, y_g])
Algorithm 1 GroupNL-based Convolutional Layer: PyTorch-like Pseudocode
3.3.2 Diversities of Data-Agnostic Nonlinear Transformation Function and Feature Maps Generation
The number of NLFs in GroupNL is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g$, which is usually larger than the fixed number of $t$ NLFs in SineFM $|\Psi_{\mathrm{sinefm}}|=t$, which act on all seed feature maps for expanding, limiting the diversity of features.
For the diversities of data-agnostic NLFs, assuming there are $64$ seed feature maps, and the number of output channels is $c_{\mathrm{out}}=256$, we can use different NLFs to generate the remaining $192$ feature maps. Assuming $t=5$, SineFM needs to generate $t\cdot c_{\mathrm{seed}}=5\cdot 64=320$ expanding features and utilizes a DWConv to align them into $192$ features. In this case, the number of NLFs is $5$. For GroupNL, the $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=192/64=3$. For $g=2$, the number of different NLFs is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 2=6$; for $g=4$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 4=12$; for $g=8$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 8=24$; and for the fully grouped $g=64$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 64=192\gg|\Psi_{\mathrm{sinefm}}|=5$. Compared with the SineFM, the diversity of NLFs is significantly improved without increasing the number of feature maps in GroupNL.
For the diversities of generated feature maps, we can visualize a given input image processing with 24-different NLFs as an example, as shown in Fig. 3. The richness of the generated feature maps is greatly improved while maintaining the similarity between different features.
Figure 3: The visualization of feature maps generated based on the diverse NLFs with 24-different hyperparameter settings. The diversity of the generated feature maps is greatly improved while maintaining the similarity.
3.3.3 Selection of NLFs, Hyperparameters and Coefficients
Choosing appropriate nonlinear transformations and selecting appropriate hyperparameters for the NLF is highly related to the performance of the GroupNL model.
Nonlinear Functions: It is worth noting that vanilla convolution operations utilize nonlinear functions for activation, which motivates us to consider the power of nonlinear transformation in feature map generation. Theoretically, all nonlinear functions can be used as NLFs in GroupNL to generate other feature maps based on the seed feature maps.
Referring to the aforementioned research about nonlinear transformations [5, 19], we select four different optimal nonlinear functions verified in these works, including the Sinusoidal function, Monomial function, Gaussian, and Laplace. Table 2 shows the specific formulas and corresponding hyperparameters of the above nonlinear functions. Furthermore, Fig. 4 illustrates the performance results of the GroupNL models with ResNet-18 backbone in four different NLFs on the CIFAR-10 dataset.
TABLE 2: Nonlinear transformation functions.
Nonlinear Transformation Formulation Hyperparameters
Sinusoidal $\mathrm{sin}(\omega(x+\phi))$ $\omega,\phi\in\mathbb{R}$
Monomial $\mathrm{sign}(x){|x|}^{\eta}$ $\eta\in\mathbb{R}$
Gaussian $e^{-{(\epsilon x)}^{2}}$ $\epsilon\in\mathbb{R}$
Laplace $\frac{\epsilon}{2}e^{-(\epsilon x)}$ $\epsilon\in\mathbb{R}$
In this case, we can observe that the GroupNL model with Sinusoidal function outperforms the other three NLFs and also has a similar performance to the vanilla CNN model. As a result, we can generate feature maps by the Sinusoidal function as the NLF in subsequent experiments.
Figure 4: Comparison of different NLFs in GroupNL ResNet-18 on the CIFAR-10 dataset.
Analysis of Sparse Seed Filters in GroupNL: As described in Section 3.3.1, we designate that the seed filters ${\bm{W}}_{\mathrm{seed}}$ in basic structure (e.g., ResNet-18) utilize the standard convolution, and the seed filters in bottleneck structure (e.g., ResNet-101) are sparse with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. The sparse seed filter experimental results are shown in Table 3. Applying sparsity to the seeded convolution of ResNet-18 can introduce an extremely low number of parameters of 0.05M and just approximately 7M FLOPs. However, performance is severely degraded. As a comparison, the ResNet-101 with sparsity can maintain the performances with the lowest number of parameters and FLOPs. Therefore, the sparse seed filters are not used in the basic structure.
TABLE 3: The ablation studies of w/ and w/o sparse seed filters and different reduction ratio $r$ in GroupNL.
Model $g$ $r$Params (M) #FLOPs (M) Top-1 Acc (%)
ResNet-18 (w/o sparse) 4 2 5.6 279 95.12
4 2.8 141 94.33
8 1.4 72 92.94
ResNet-18 (w/ sparse) 4 2 0.05 6.9 89.66
4 0.05 7.0 87.17
8 0.05 7.1 87.67
ResNet-101 (w/o sparse) 4 2 29.4 1790 95.57
4 23.9 1472 95.21
8 21.2 1312 95.06
16 19.8 1233 93.73
ResNet-101 (w/ sparse) 4 2 18.5 1159 95.51
4 18.5 1159 95.40
8 18.5 1159 95.10
16 18.5 1159 93.83
Hyperparameters and Coefficients of GroupNL: We designate that the hyperparameters of NLFs are randomly initialized and non-learnable during model training. This trick reduces the number of gradients for optimization in training while preventing the NLFs lose the data-agnostic properties. The Sinusoidal function has two hyperparameters $\omega$ and $\phi$, we randomly choose them via $\tt{torch.rand}$ between $[1,2]$ and $[1,5]$, respectively.
In addition, we also explore the reduction ratio $r=c_{\mathrm{out}}/c_{\mathrm{seed}}$ and the number of splitting groups $g$. Ablation studies of different configurations on $r$ and $g$ are shown in Table 3 and Table 4, respectively. As shown in Table 3, the model’s accuracy decreases as the reduction ratio increases. As shown in Table 4, with the fixing $r$, the change in $g$ results in model performance improvement in proper $g$.
Besides, for the ResNet-101, the number of parameters and FLOPs are very similar for different $r$ settings. This is due to the bottleneck structure in ResNet-101 using the sparse seed filters with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. For example, assuming a $k^{2}$ convolution with same input and output channels $c_{\mathrm{in}}=c_{\mathrm{out}}=512$ and feature maps size $wh$ in the bottleneck. In the case of $r=2$, the number of parameters is $\mathcal{P}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}}{256}=512k^{2}$, and the FLOPs is $\mathcal{F}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}wh}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}wh}{256}=512k^{2}wh$. In the cases of $r=\{4,16,32,\cdots,256,512\}$, the results are also $\mathcal{P}_{r}=512k^{2}$ and $\mathcal{F}_{r}=512k^{2}wh$ because of the fraction $c_{\mathrm{seed}}$ are reduced in the case of using the greatest common divisor grouping seed filters.
For the $g$, on the other hand, because the parameters and computations depend on the seed filters, different $g$ will not affect the FLOPs and the number of parameters.
TABLE 4: The ablation studies of different splitting groups $g$ in GroupNL CNN models.
Model $r$ $g$Params (M) #FLOPs (M) Top-1 Acc (%)
ResNet-18 2 4 5.6 279 95.12
8 5.6 279 94.91
16 5.6 279 94.84
4 4 2.8 141 94.33
8 2.8 141 94.38
16 2.8 141 94.15
ResNet-101 2 4 18.5 1159 95.51
8 18.5 1159 95.35
16 18.5 1159 95.66
4 4 18.5 1159 95.40
8 18.5 1159 95.44
16 18.5 1159 95.25
3.4 Resource and Training Acceleration Analysis
For cloud-assisted CNN models, the multi-GPUs training on cloud servers is mainly based on data parallelism. In practical multi-GPU training, the bandwidth of inter-GPU memory data transfer is an important constraint on the training speed compared to the powerful and fast computation kernels. In the classic data parallelism, e.g., $\tt{torch.nn.DataParallel}$, the data exchange between different GPUs in forward/backward of the model generally in three stages: (i) Gather outputs (or logits) and compute loss gradients on host GPU; (ii) Scatter gradients to multi-GPUs; (iii) Reduce gradients to host GPUs. For the distributed data parallelism, e.g., $\tt{torch.nn.parallel.DistributedDataParallel}$, only one reduction of gradients between GPUs in each backward.
So, the number of gradients can be further optimized in terms of inter-GPU scatters and reductions in different data parallelism. Assuming each trainable parameter in a CNN model maintains a gradient, the number of gradients $\mathcal{G}$ corresponds to the number of trainable parameters $\mathcal{P}$.
At the same time, designing libraries-agnostic and task-specific CNN methods is the minimal change in cloud-assisted architecture for generalization. Based on this premise, we analyze the theoretical gradient exchange in multi-GPUs memory of vanilla CNN, GhostNet, SineFM, and GroupNL in training.
Ignoring the bias, for a vanilla convolutional layer, the number of parameters and gradients can be described as follows:
$\begin{array}[]{l}{\mathcal{P}_{\mathrm{vanilla\ cnn}}=\mathcal{G}}_{\mathrm{vanilla\ cnn}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\ .\end{array}$ (13)
For a GhostNet layer, the number of parameters and gradients from the seed filters and $d\times d$ DWConv are:
${\mathcal{P}_{\mathrm{ghostnet}}=\mathcal{G}}_{\mathrm{ghostnet}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+c_{\mathrm{gen}}d^{2}$ (14)
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!-\!\frac{1}{r})d^{2}\ ,$
and the number of parameters and gradients for the SineFM layer from the seed filters, $1\times 1$ DWConv, and batch normalization are:
${\mathcal{P}_{\mathrm{sinefm}}=\mathcal{G}}_{\mathrm{sinefm}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+tc_{\mathrm{gen}}\cdot 1^{2}+2tc_{\mathrm{seed}}$ (15)
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+tc_{\mathrm{out}}(1\!-\!\frac{1}{r})+2tc_{\mathrm{out}}/r$
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!+\!\frac{1}{r})t\ .$
The proposed GroupNL layer only contains the learnable standard seed filters and learnable sparse seed filters, and the number of parameters and gradients are shown in Eq. 16 and Eq. 17, respectively:
[…]
Noticeably, the relation is $\mathcal{G}_{\mathrm{groupnl}}^{\mathrm{sparse}}\leq\mathcal{G}_{\mathrm{groupnl}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r<\mathrm{min}\{{\mathcal{G}}_{\mathrm{cnn}},{\mathcal{G}}_{\mathrm{ghostnet}},{\mathcal{G}}_{\mathrm{sinefm}}\}$, i.e., the number of gradients for optimization in GroupNL is minimal compared to the rest CNN design methods of ${\mathcal{G}}_{\mathrm{cnn}}$, ${\mathcal{G}}_{\mathrm{ghostnet}}$, and ${\mathcal{G}}_{\mathrm{sinefm}}$. The sparse GroupNL also significantly reduces the gradient exchange by a factor of $\frac{1}{r\cdot\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{out}}/r)}$ compared to the ${\mathcal{G}}_{\mathrm{cnn}}$. As the number of GPUs scales, this scatter/reduction gap compared to vanilla CNNs also grows larger thus accelerating the training.
It is notable that, it seems that GhostNet and SineFM also have a small number of gradients with ${{\mathcal{G}}_{\mathrm{ghostnet}}<\mathcal{G}}_{\mathrm{cnn}}$ and ${\mathcal{G}}_{\mathrm{sinefm}}<{\mathcal{G}}_{\mathrm{cnn}}$, can also accelerate training in theory. However, the evaluation in Table I shows that the GhostNet and SineFM are even slower than the vanilla CNN. For this phenomenon, we consider this to be due to the introduction of more computational nodes involved in the backward updating, such as DWConv and Norm, which leads to slower training instead. Our proposed GroupNL, on the other hand, does not introduce new computational nodes through in-place operations such as $\tt{torch.cat}$ and $\tt{torch.split}$, and copy operations such as $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, thus ensures acceleration of training.
3.5 Discussion
As the aforementioned analysis, compared with the GhostNet [18] and SineFM [19], the proposed GroupNL-based CNN design method has two highlights:
(i) More diverse feature maps. Given the seed filters and the number of output channels of the layer, by grouping the seed filters, more different hyperparameter settings of data-agnostic NLFs can be used to generate diverse features.
(ii) More training efficient. GroupNL generates the number of feature maps without extra computational nodes like DWConv and Norm, effectively reducing the number of gradients exchanging in the multi-GPUs training.
4 Experiments
This section introduces our experimental setup, including the datasets, baselines, and implementation details. Then, we empirically compare accuracy, network traffic, FLOPs, and training speed on multiple benchmarks.
4.1 Experimental Setup
4.1.1 Datasets
We evaluate the effectiveness of the proposed method using five widely used datasets. CIFAR-10, GTSRB, and ImageNet-1K are standard datasets. CIFAR-10-C and Icons-50 are used as corrupted datasets. Among them, noise is introduced in the CIFAR-10-C dataset, and the training data and test data styles of Icons-50 are quite different.
CIFAR-10 [35] is a multi-class natural object dataset for image classification. It consists of 50,000 training images and 10,000 test images in 10 categories, each with a resolution of $32\times 32$ pixels.
GTSRB [36] is a German traffic sign dataset for traffic sign recognition. It comprises 39,209 training images and 12,630 test images from 43 categories. The images in GTSRB have varying light conditions and rich backgrounds. We set the resolution of all images to $32\times 32$ pixels.
CIFAR-10-C [37] is a test dataset after using synthetic common perturbations and noise corruptions on the CIFAR-10 test set. It consists of 10,000 test images of 19 types of damage in 4 categories. The resolution of each image is $32\times 32$ pixels.
Icons-50 [37] consists of 10,000 images in 50 categories collected from different companies. We set the resolution of all images to $32\times 32$ pixels. For training, the data of one company is retained as test data, and the data of the rest of the companies is used as training data.
Please refer to [5] for more explanation of the CIFAR-10-C and Icons-50 datasets.
ImageNet [38] is a large-scale classification dataset to verify computer vision performance. ImageNet-1K is a subset of the ILSVRC ImageNet, which contains 1.28 billion training images and 50K validation images from 1K different classes. The resolution of each image is $224\times 224$ pixels.
4.1.2 Baseline CNNs
We compare the following nonlinear transformation and feature maps generated-based methods in CNNs.
MonoCNN [5]: In MonoCNN, the convolutional layer is specified with only a few seed filter parameters updated during model training. The remaining filters in this layer are generated based on the seed filters and monomial function. The hyperparameters of the monomial function are randomly generated and non-learnable in training.
GhostNet [18]: In GhostNet, a few filters are first used to generate some feature maps. Then, utilizing $d\times d$ DWConv operations to generate the remaining feature maps based on these generated feature maps.
SineFM [19]: In the SineFM-based CNN model, a small number of seed convolutional filters are first specified in the convolutional layer. Then, seed feature maps are generated based on the seed filters and input data. Subsequently, a set of expanding feature maps is generated based on the seed feature map, $t$-sinusoidal functions, and a $t$-sets of BatchNorm. Finally, utilizing the $1\times 1$ DWConv to align the number of expanding features to the generated features.
To ensure a comprehensive comparison, all baseline methods are based on the following CNN models: ResNet-18/-34/-101 [2] and VGG11 [1]. To make the effect of training acceleration more clear, we evaluate the training speed of the larger model, i.e., the ResNet-101.
4.1.3 Implementation Details
The implementation details include training platforms & libraries and experimental settings.
Training Platforms and Libraries: The experiments are performed on the servers with NVIDIA RTX 2080Ti 11GB and RTX 4090 GPUs 24GB. We implement the models in PyTorch 1.12 for the NVIDIA platforms with CUDA 11.1.
Hyperparameters Settings of Baseline CNNs: For the hyperparameters setting, we follow the suggestion from the original papers [5, 18, 19]. The specific description of model hyperparameters is shown in Table 5.
(i) MonoCNN [5] uses the monomial function as a filter generating function and samples a random and continuous exponent $\eta$ between $[1,7]$, and the exp_factor is $\{2,4,8,16\}$;
(ii) GhostNet [18] uses $3\times 3$ grouped convolution with $d=3$ as the cheap operation for generating the feature maps and sets the reduction ratio to 2;
(iii) SineFM [19] uses a sinusoidal function as the feature map generator function with random hyperparameters, the real numbers of period $\omega$ and shift $\phi$, which ranges are between $[1,2]$ and $[1,5]$, respectively. The number of the NLFs with different hyperparameter settings in SineFM is $t=5$, which means five times expanding the seed features as described in Eq. 7.
Experimental Settings: For the settings of training in CIFAR-10/GTSRB/Icons-50, the learning rate is 0.1, and the batch size is 128, with the SGD optimizer and cosine annealing scheduler. For the training in ImageNet-1K, the learning rate is 0.001, and the batch size is 1024, with the Adam optimizer and cosine annealing scheduler.
In addition, the training evaluation is compared for the larger model, i.e., ResNet-101 in different parallel technologies in PyTorch library API [39], including the Data Parallelism (DP) with $\tt{torch.nn.DataParallel}$, the Distributed Data Parallelism (DDP) [40] with $\tt{torch.nn.parallel.DistributedDataParallel}$, and the Automatic Mixed Precision (AMP) [41] with $\tt{torch.cuda.amp}$. We also follow the linear scaling rule [42] to scale the number of parallelism in training, which is scaling the learning rate by the batch size. In addition, we also set up the same scale of seed feature map numbers in GhostNet, SineFM, and GroupNL with $r=2$.
TABLE 5: Hyperparameter setting of CNNs methods.
Method Hyperparameters Settings
Reduction Coefficient & Random Range
MonoCNN $\mathrm{exp\_factor}=\{2,4,8,16\}$ $\eta\in[1,7]$
GhostNet $r=2$ $d=3$
SineFM $r=2$ $t=5,\quad\omega\in[1,2],\quad\phi\in[1,5]$
GroupNL $r=2$ $g=4,\quad\omega\in[1,2],\quad\phi\in[1,5]$
4.2 Experimental Results
4.2.1 Reduction of Network Traffic and FLOPs
Figure 5: Number of learnable parameters for different models with different methods.
Figure 6: FLOPs for different models with different methods.
Since the non-learnable hyperparameters of NLFs are randomly initialized and remain unchanged, these parameters can be saved and reproduced with a random number. Therefore, the number of learnable parameters in the CNN model can be approximated to the number of model parameters that need to be transmitted when deploying the model on cloud-assisted IoT devices.
Fig. 5 and Fig. 6 show the number of learnable parameters and FLOPs of different CNN models for different methods. Without using sparse technology, the number of learnable parameters of GhostNet, SineFM, and GroupNL is similar, which is much less than the number of learnable parameters of vanilla CNN. As shown in Fig. 5, When based on VGG11, ResNet-18, and ResNet-34 architectures, GhostNet, SineFM, and GroupNL have a similar number of learnable parameters. When based on ResNet-101 architecture, GroupNL has the least amount of learnable parameters due to the user of sparse technology. As shown in Fig. 6, in terms of FLOPs, there are similar findings as in the amount of learnable parameters.
4.2.2 Robustness
This section evaluates the robustness of the GroupNL-based CNN on three types of datasets, including standard datasets (i.e., CIFAR-10 and GTSRB), the corrupted dataset (i.e., CIFAR-10-C), and data under different styles (i.e., Icons50).
1) Standard Data.
TABLE 6: Comparison of image classification results on CIFAR-10 and GTSRB datasets.
Methods CIFAR-10 Top-1 Acc (%) GTSRB Top-1 Acc (%)
ResNet-18 ResNet-34 ResNet-101 VGG11 ResNet-18
Vanilla CNN 95.04 95.24 95.48 91.89 97.89
MonoCNN 92.05 92.26 94.97 88.84 97.66
GhostNet 93.36 93.90 95.04 92.65 95.76
SineFM 94.36 94.77 95.06 91.27 98.18
GroupNL 95.12 95.45 95.51 91.62 98.31
Table 6 shows the comparison of image classification results of different CNN models on CIFAR-10 and GTSRB datasets under different methods. We observe that:
(i) The accuracy of the GroupNL-based CNN exceeds that of all alternatives to the standard CNN model. As shown, when using ResNet-34 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.68% higher than SineFM’s, 1.55% higher than GhostNet’s, and 3.19% higher than MonoCNN’s. When using VGG11 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.35% higher than SineFM’s. The main reason is that GroupNL generates diversified feature maps using the NLFs with different hyperparameter configurations. In contrast, SineFM generates feature maps using a single NLF and is unable to generate diverse feature maps. GhostNet first generates some feature maps using standard convolution operations and then uses cheap linear operations to generate other feature maps based on these generated feature maps, which limits the diversity of the feature maps. Benefiting from its ability to generate diverse feature maps, GroupNL achieves the highest accuracy among many alternatives.
(ii) The GroupNL-based CNN models outperform the vanilla CNN models in most cases. As shown in Table 6, GroupNL performs slightly worse than the vanilla CNN only when using VGG11 as the backbone. When using ResNet-18/-34/-101, the accuracy of GroupNL exceeds that of the vanilla CNN. This is mainly due to the ability of GroupNL to generate diverse feature maps.
2) Corrupted Data.
TABLE 7: Robustness to common observable corruptions. Our results are highlighted with shading.
Method Noise Blur Weather Digital mean
Vanilla CNN $52.60$ $72.78$ $84.55$ $80.70$ $72.58$
MonoCNN $56.36$ $63.24$ $77.79$ $75.17$ $68.14$
GhostNet $59.26$ $72.57$ $82.99$ $79.95$ $73.69$
SineFM $53.77$ $68.60$ $84.63$ $80.94$ $71.99$
GroupNL $57.55$ $72.96$ $85.25$ $80.68$ $74.11$
Method Noise Blur Weather Digital mean
Vanilla CNN $53.72$ $71.70$ $85.08$ $82.27$ $73.19$
MonoCNN $57.16$ $63.50$ $78.78$ $75.81$ $68.81$
GhostNet $60.30$ $74.57$ $84.10$ $80.71$ $74.92$
SineFM $59.20$ $71.72$ $86.11$ $81.83$ $74.72$
GroupNL $59.84$ $73.64$ $86.01$ $81.26$ $75.19$
Method Noise Blur Weather Digital mean
Vanilla CNN $53.97$ $73.05$ $86.02$ $81.90$ $74.78$
MonoCNN $49.03$ $74.36$ $86.60$ $82.00$ $74.26$
GhostNet $55.03$ $74.00$ $86.13$ $82.45$ $74.40$
SineFM $58.33$ $72.08$ $85.08$ $81.44$ $74.23$
GroupNL $57.90$ $71.22$ $86.45$ $81.17$ $75.04$
Method Noise Blur Weather Digital mean
Vanilla CNN $65.37$ $75.45$ $82.06$ $80.48$ $75.84$
MonoCNN $42.74$ $68.73$ $78.08$ $76.58$ $66.53$
GhostNet $52.91$ $70.28$ $80.05$ $78.65$ $70.47$
SineFM $62.31$ $72.46$ $82.46$ $82.61$ $74.96$
GroupNL $64.10$ $77.24$ $83.46$ $83.87$ $77.17$
Table 7 shows the performance comparison of different models for each method on the corrupted data. We observe that:
The GroupNL-based CNN models achieve the highest accuracy. As shown in Table 7, when ResNet-34 is used as the backbone, the accuracy of GroupNL is 0.47% higher than that of SineFM, 6.38% higher than that of MonoCNN, 0.27% higher than that of GhostNet, and even 2% higher than the accuracy of standard ResNet-34. When VGG11 is used as the backbone, the accuracy of GroupNL is 2.21% higher than that of SineFM, 10.64% higher than that of MonoCNN, 6.7% higher than that of GhostNet, and 1.33% higher than the accuracy of standard VGG11. There are three reasons for this: (i) GroupNL sets some model parameters to be non-learnable, so that the model parameters are not completely obtained through training by inputting training data, thus making the model have stronger generalization ability. (ii) The rules of the NLF can regularize the model and improve the generalization ability of the model. (iii) NLFs equipped with different hyperparameters can generate diverse feature maps and improve the knowledge expression ability of the model. Therefore, GroupNL achieves the highest accuracy.
The model’s ability to generate diverse feature maps has a great impact on the robustness of the model. As shown, with ResNet-18 as the backbone, GroupNL’s accuracy is 2.12% higher than SineFM’s. With ResNet-34 as the backbone, GroupNL’s accuracy is 0.47% higher than SineFM’s. With ResNet-101 as the backbone, GroupNL’s accuracy is 0.81% higher than SineFM’s. With the same number of parameters and computational complexity, the difference between GroupNL and SineFM is that the former can generate a variety of feature maps. This shows that increasing the diversity of feature maps can effectively improve the robustness of the CNN model.
TABLE 8: Robustness to different styles. Our results are highlighted with shading.
Method Apple Facebook Google Samsung mean
Vanilla CNN $93.66$ $87.64$ $83.20$ $80.28$ $86.19$
MonoCNN $93.02$ $85.37$ $81.60$ $78.59$ $84.65$
GhostNet $92.07$ $85.09$ $78.98$ $77.61$ $83.43$
SineFM $94.84$ $88.72$ $84.96$ $82.27$ $87.70$
GroupNL $94.52$ $91.05$ $86.27$ $84.36$ $89.05$
Method Apple Facebook Google Samsung mean
Vanilla CNN $92.77$ $86.26$ $83.81$ $82.44$ $86.32$
MonoCNN $91.46$ $85.88$ $81.55$ $79.23$ $84.53$
GhostNet $89.87$ $84.70$ $79.61$ $73.76$ $81.98$
SineFM $94.74$ $88.72$ $84.88$ $81.46$ $87.45$
GroupNL $95.19$ $90.98$ $85.93$ $84.36$ $89.12$
Method Apple Facebook Google Samsung mean
Vanilla CNN $93.64$ $87.70$ $83.64$ $83.13$ $87.03$
MonoCNN $94.84$ $87.64$ $83.62$ $82.71$ $87.20$
GhostNet $93.50$ $88.53$ $83.62$ $84.50$ $87.54$
SineFM $90.25$ $83.65$ $79.69$ $74.67$ $82.06$
GroupNL $93.98$ $90.15$ $85.51$ $83.89$ $88.38$
Method Apple Facebook Google Samsung mean
Vanilla CNN $92.54$ $86.68$ $83.67$ $80.34$ $85.81$
MonoCNN $89.55$ $81.10$ $79.82$ $75.41$ $81.47$
GhostNet $91.40$ $88.15$ $80.71$ $79.33$ $84.90$
SineFM $94.74$ $89.87$ $85.59$ $82.17$ $88.09$
GroupNL $94.23$ $91.05$ $85.30$ $82.13$ $88.18$
3) Data under different styles: We obtain conclusions similar to those for the corrupted data; the GroupNL-based CNN models achieve the highest accuracy under all backbones. As shown in Table 8, when ResNet-18 is used as the backbone, the accuracy of GroupNL is 1.35% higher than that of SineFM, 4.4% higher than that of MonoCNN, 5.62% higher than that of GhostNet, and even 2.86% higher than the accuracy of standard ResNet-18. when ResNet-101 is used as the backbone, the accuracy of GroupNL is 6.3% higher than that of SineFM, 1.18% higher than that of MonoCNN, 0.84% higher than that of GhostNet, and even 1.35% higher than the accuracy of standard ResNet-18. The reason for this is the same as that of our corrupted data; that is, (i) GroupNL makes the model parameters not completely updated by the input training data; (ii) the nonlinear transformation function in GroupNL regularizes the model; (iii) GroupNL can generate diverse feature maps. Therefore, GroupNL has the highest accuracy.
4.2.3 Training Acceleration
TABLE 9: Comparison of CIFAR-10 Training Time in Different CNN methods in RTX 2080Ti GPUs.
Tech. Parallelism No. of GPUs LR BS Method Top-1 Acc (%) Training Time Accelerated Ratio
Data Parallel $\times$1 2-GPUs 0.1 64$\cdot$2 Vanilla CNN 95.48 9.3 Hours -
MonoCNN 94.97 9.2 Hours +0%
GhostNet 95.04 12.7 Hours -37%
SineFM 95.06 15.4 Hours -66%
GroupNL 95.51 8.3 Hours +11%
We verify the training acceleration of each CNN method based on ResNet-101. We observe that:
(i) Compared to vanilla ResNet-101, GroupNL-based ResNet-101 achieves 11% speedup while achieving the higher accuracy, as shown in Table 9. This is because the GroupNL method uses the NLFs to generate feature maps, reducing the number of operations based on sliding windows to generate feature maps. The FLOPs of generating feature maps based on nonlinear transformation are significantly less than those of generating feature maps based on sliding windows (as shown in Fig. 6). Moreover, since only two GPUs are used, the GroupNL mainly accelerates model training by reducing model calculation time.
(ii) Although the GhostNet and SineFM have fewer parameters and computational complexity than vanilla CNN, the training time is longer, as shown in Table 9. As we analyzed in Section 2.1, although cheap linear transformations and normalizations require less computation, they are slow to compute and optimize. Incorporating these operations into the model will significantly slow down the model’s training speed.
We then further compare the speedup of GroupNL and vanilla CNN with different numbers of GPUs under Data Parallel (DP) and Distributed Data Parallel (DDP).
1) Training Evaluation in DP. For the multi-GPUs evaluation using DP technology, Table 10 and Fig. 7 show the experimental settings and training results, respectively. We observe that:
TABLE 10: Training Settings in DP Evaluation.
Platform Parallelism No. of GPUs LR No. of BS
RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
$\times$1 2-GPUs 0.1 64$\cdot$2
$\times$2 2-GPUs 0.2 128$\cdot$2
$\times$3 3-GPUs 0.3 128$\cdot$3
$\times$4 4-GPUs 0.4 128$\cdot$4
With more GPUs, the GroupNL has a larger training speed-up ratio. As shown in Fig. 7, when the number of GPUs is 4, GroupNL has the largest speed-up ratio compared to the vanilla CNN. This is because when the number of GPUs is large in DP, the parameter transmission time between multiple GPUs will increase significantly when training vanilla CNN. In contrast, the non-trainable parameters in the GroupNL do not need to be transmitted and are less affected by the parameter transmission between multiple GPUs. Therefore, when a large number of GPUs are used, GroupNL has a higher training speedup ratio.
Figure 7: CIFAR-10 Training Acceleration Evaluation with Data Parallelism in RTX 2080Ti GPUs.
2) Training Evaluation in DDP. For the multi-GPUs evaluation with DDP technology, Table 11 and Fig. 8 show the experimental settings and evaluated results of DDP training with AMP, respectively. We observe that: Since DDP uses Ring-AllReduce to reduce parameter transmission between multiple GPUs, the GroupNL reduces training time compared to the vanilla CNN mainly because it reduces the computational workload of the CNN model. This shows that the GroupNL can effectively speed up the training of the model by using nonlinear transformation to generate feature maps to reduce the use of sliding windows to generate feature maps, and also makes the GroupNL-based model more suitable for deployment on IoT devices.
TABLE 11: Training Settings in DDP Evaluation.
Platform Parallelism No. of GPUs LR No. of BS
RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
$\times$2 2-GPUs 0.2 128$\cdot$2
$\times$4 2-GPUs 0.4 256$\cdot$2
$\times$8 4-GPUs 0.8 256$\cdot$4
RTX 4090 $\times$8 8-GPUs 0.001 128$\cdot$8
Figure 8: CIFAR-10 Training Acceleration Evaluation with DDP and AMP in RTX 2080Ti GPUs.
Figure 9: ImageNet-1K Training Acceleration Evaluation with DDP (Parallelism ×8) in RTX 4090 GPUs.
We also show the training speedup results for ImageNet-1K training of 8 GPUs. The GroupNL reduces training time by about 53% while achieving similar accuracy to the vanilla CNN. As shown in Fig 9. The $\times$8 DDP training vanilla CNN for 10 epochs is 3.32 hours with a Top-1 accuracy of 52.60%, and GroupNL is 1.87 hours with a Top-1 accuracy of 52.65%. For the $\times$8 DDP training vanilla CNN for 300 epochs is 151.93 hours with a Top-1 accuracy of 81.04%, and GroupNL is 72.11 hours with a Top-1 accuracy of 80.16%.
In general, GroupNL can accelerate model training because: (i) the non-trainable parameters contained in GroupNL do not need to be transferred between multiple GPUs, reducing the transmission time during model training; (ii) nonlinear transformation operations replace most operations based on sliding window feature map generation, reducing the calculation time during model training. GroupNL optimizes the model acceleration in terms of parameter transmission and calculation time. Therefore, it can accelerate model training.
5 Conclusion and Future Work
This paper presented a robust and low-resource CNN design method, GroupNL. In GroupNL, seed feature maps are grouped and nonlinear transformation functions configured with different hyperparameters are used to generate diverse feature maps on demand to reduce the number of parameters and floating-point operations (FLOPs). Experimental results show that GroupNL can effectively improve the robustness of the CNN model on multiple tasks while reducing the number of model parameters and FLOPs.

### v2

@@
 Abstract
-It has become mainstream to deploy Convolutional Neural Network (CNN) models on ubiquitous Internet of Things (IoT) devices with the help of the cloud to provide users with a variety of high-quality services.
-Most existing methods have two limitations: (i) low robustness in handling corrupted image data collected by IoT devices; and (ii) high consumption of computational and transmission resources.
-To this end, we propose the Grouped NonLinear transformation generation method (GroupNL), which generates diversified feature maps by utilizing data-agnostic Nonlinear Transformation Functions (NLFs) to improve the robustness of the CNN model.
-Specifically, partial convolution filters are designated as seed filters in a convolutional layer, and a small set of feature maps, i.e., seed feature maps, are first generated based on vanilla convolution operation.
-Then, we split seed feature maps into several groups, each with a set of different NLFs, to generate corresponding diverse feature maps with in-place nonlinear processing.
-Moreover, GroupNL effectively reduces the parameter transmission between multiple nodes during model training by setting the hyperparameters of NLFs to random initialization and not updating them during model training, and reduces the computing resources by using NLFs to generate feature maps instead of most feature maps generated based on sliding windows.
-Experimental results on CIFAR-10, GTSRB, CIFAR-10-C, Icons50, and ImageNet-1K datasets in NVIDIA RTX GPU platforms show that the proposed GroupNL outperforms other state-of-the-art methods in model robust and training acceleration.
-Specifically, on the Icons-50 dataset, the accuracy of GroupNL-ResNet-18 achieves approximately 2.86% higher than the vanilla ResNet-18.
-GroupNL improves training speed by about 53% compared to vanilla CNN when trained on a cluster of 8 NVIDIA RTX 4090 GPUs on the ImageNet-1K dataset.
+Deploying Convolutional Neural Network (CNN) models on ubiquitous Internet of Things (IoT) devices in a cloud-assisted manner to provide users with a variety of high-quality services has become mainstream.
+Most existing studies speed up model cloud training/on-device inference by reducing the number of convolution (Conv) parameters and floating-point operations (FLOPs).
+However, they usually employ two or more lightweight operations (e.g., depthwise Conv, $1\times 1$ cheap Conv) to replace a Conv, which can still affect the model’s speedup even with fewer parameters and FLOPs.
+To this end, we propose the Grouped NonLinear transformation generation method (GroupNL), leveraging data-agnostic, hyperparameters-fixed, and lightweight Nonlinear Transformation Functions (NLFs) to generate diversified feature maps on demand via grouping, thereby reducing resource consumption while improving the robustness of CNNs.
+First, in a GroupNL Conv layer, a small set of feature maps, i.e., seed feature maps, are generated based on the seed Conv operation.
+Then, we split seed feature maps into several groups, each with a set of different NLFs, to generate the required number of diversified feature maps with tensor manipulation operators and nonlinear processing in a lightweight manner without additional Conv operations.
+We further introduce a sparse GroupNL Conv to speed up by reasonably designing the seed Conv groups between the number of input channels and seed feature maps.
+Experiments conducted on benchmarks and on-device resource measurements demonstrate that the GroupNL Conv is an impressive alternative to Conv layers in baseline models.
+Specifically, on Icons-50 dataset, the accuracy of GroupNL-ResNet-18 is 2.86% higher than ResNet-18; on ImageNet-C dataset, the accuracy of GroupNL-EfficientNet-ES achieves about 1.1% higher than EfficientNet-ES.
+In addition, we verified the efficiency of GroupNL-based models in terms of cloud training and on-device inference.
 Index Terms:
@@
 1 Introduction
-Internet of Things (IoT) Analytics predicts the number of connected IoT devices worldwide will increase from 16.7 billion in 2023 to 29 billion in 2027111https://iot-analytics.com/number-connected-iot-devices/.
-It is possible to provide convenient services with the help of connected and ubiquitous IoT devices.
-Additionally, given the excellent performance of convolutional neural networks (CNN) [1, 2, 3, 4] in computer vision applications, it has become mainstream to deploy CNN models on IoT devices to provide people with various convenient and high-quality services, such as image recognition and monitoring services [5, 6, 7, 8, 9].
+Internet of Things (IoT) Analytics predicts that the number of connected IoT devices worldwide will increase from 16.7 billion in 2023 to 29 billion in 2027111https://iot-analytics.com/number-connected-iot-devices/.
+It is possible to provide convenient services using connected, ubiquitous IoT devices.
+Due to the limited computing and storage resources of IoT devices, using the cloud server to train Convolutional Neural Netowrks (CNNs) [35, 13, 18, 6, 16, 34, 17, 8, 36, 7, 37] and then sending them to IoT devices for deployment and service provision (such as image recognition and monitoring services [3, 45, 44, 33, 39]) has become mainstream, as shown in Fig.
+1.
 Figure 1: Overview of cloud-assisted architecture.
-IoT devices such as smart cameras and monitors usually work outdoors and often need to process data affected by the natural environment (such as fog, rain, snow, etc.).
-Vanilla CNN models have low robustness when processing such environmentally affected data.
-In addition, IoT devices have limited computing and storage resources.
-Using cloud servers to train CNN models and then sending them to IoT devices for deployment and service provision has become mainstream, as shown in Fig.
-1.
-However, a cloud server usually corresponds to thousands of IoT devices, and frequent deployment and updating of vanilla CNN models will occupy a lot of bandwidth resources.
-In addition, the vanilla CNN generates feature maps through sliding windows, which consumes a lot of floating-point operations (FLOPs).
-Therefore, to provide high-quality services on IoT devices, it is urgent to design a CNN model that is highly robust, parameter transmission-friendly and has low computing resource requirements.
-This paper aims to propose a CNN design method that increases the diversity of feature maps to improve model robustness without increasing model complexity while ensuring parameter transmission friendliness when deploying and updating models for cloud-assisted IoT devices.
-To this end, first, this paper proposes the Grouped NonLinear transformation method, GroupNL.
-In the GroupNL-based CNN model, partial trainable filters are configured within each convolution layer, which is referred to as seed filters.
+Given the limited resources of IoT devices and users’ demand for fast service response, many methods have been proposed to reduce the resource consumption of CNN models and speed up model training/inference by reducing the number of convolutional (Conv) parameters and floating-point operations (FLOPs) during feature map generation.
+For example, GhostNet [11] uses vanilla Conv and lightweight Conv (e.g., depthwise Conv [16]) to generate feature maps.
+SineFM [27] combines vanilla Conv, nonlinear transformation functions (NLFs) with fixed hyperparameters, and $1\!\times\!1$ cheap Conv [11] to generate feature maps.
+These methods can reduce the number of model parameters and FLOPs, but the lightweight operations (depthwise Conv and cheap Conv) introduced affect the model’s cloud training/on-device inference speed.
+In addition, fixing hyperparameters of NLFs prevents model parameter updates from being completely dependent on the training data, regularizing the model to some extent and improving its robustness, while effectively enhancing the robustness of lightweight CNN models to environmental factors is a task of significant value that remains to be explored.
+To this end, this paper aims to propose a CNN design method that improves model robustness by increasing feature map diversity, while reducing model complexity and avoiding the introduction of lightweight operations such as depthwise Conv and cheap Conv to achieve better training/inference speed.
+First, this paper proposes the Grouped NonLinear transformation method, GroupNL.
+In the GroupNL CNN, partially trainable filters are configured within each Conv layer, which are referred to as seed filters or seed Conv.
 The feature maps produced via seed filters are called seed feature maps.
 Subsequently, we split the seed feature maps into many small groups.
-The remaining feature maps can be generated with diverse nonlinear transformation functions (NLFs) configured with different hyperparameters for each group of seed feature maps.
-The seed feature maps and generated feature maps are concatenated as the output of the convolutional layer.
-GroupNL has four advantages: (i) By grouping seed feature maps, GroupNL can generate diversified feature maps using NLFs with different hyperparameters without increasing the complexity of the model, effectively improving the robustness of the model.
-(ii) The hyperparameters of NLFs are randomly initialized and remain unchanged during model training.
-They can be saved and reproduced using random numbers so that after the cloud trains the model, only a small amount of learning parameters and random numbers need to be transmitted.
-Compared with transmitting the complete model, the amount of parameters transmitted to the IoT device can be effectively reduced.
-In addition, when the GroupNL-based CNN model employing parallel training technologies [10, 11, 12] like data parallelism (DP) and distributed data parallelism (DDP) to accelerate model training, only a small number of gradients in seed filters have to exchange between multi-GPUs for scatter and gather.
-This process reduces the intermediate data exchange between different GPUs.
-Compared with scattering and gathering all gradients in the vanilla CNN model, only exchanging a small number of gradients from seed filters in multi-GPUs memory dramatically accelerates the training speed.
-(iii) GroupNL includes generating feature maps based on sliding windows and nonlinear transformations.
-Compared with the vanilla CNN that generates feature maps completely based on sliding windows, it can effectively reduce the number of FLOPs required to run the CNN model.
-(iv) GroupNL is simple and easy to use, and the convolution modules it generates can directly replace the convolution layers in the existing CNN architecture.
-Then, for larger CNN models, such as ResNet-101, this paper introduces sparse technology, that is, apply group convolution to generate seed feature maps, which further reduces the number of learnable parameters and FLOPs of the GroupNL-based model while ensuring model performance.
-This paper also analyzes four types of NLFs, i.e., Sinusoidal, Monomial, Gaussian, and Laplace functions.
-It verifies on multiple datasets that the feature maps generated based on the Sinusoidal function have the best performance.
-Finally, this paper conducts extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, Icons-50, and ImageNet-1K datasets, and experimental results demonstrate that our proposed GroupNL-based CNN models achieve strong performance in robustness and resource-saving.
-For example, the accuracy of GroupNL ResNet-34 on CIFAR-10-C is about 2.0% higher than that of vanilla ResNet-34, and the accuracy of GroupNL ResNet-18 on Icons-50 is about 2.86% higher than that of vanilla ResNet-18.
-We also verify the training acceleration in CIFAR-10 and ImageNet-1K datasets on the NVIDIA RTX 2080Ti/RTX 4090 multi-GPUs platforms.
-Particularly, with the distributed data parallel training in ImageNet-1K, the GroupNL ResNet-101 achieves an appropriately 53% speedup compared to the vanilla ResNet-101 while maintaining accuracy.
+The remaining feature maps can be generated using diverse NLFs, each configured with different hyperparameters for each group of seed feature maps.
+The hyperparameters of NLFs are fixed, ensuring transmission friendliness when cloud-assisted.
+The seed feature maps and the generated feature maps are concatenated into the Conv layer’s output.
+The generation process of GroupNL consists entirely of tensor manipulation operators and nonlinear transformation operations.
+Compared with GhostNet and SineFM, there is no additional cheap Conv and Batch Normalization.
+Therefore, GroupNL can achieve better model training/inference acceleration effects.
+Then, this paper designs the sparse GroupNL Conv by reasonably setting the relationship between the seed feature map ratio and the number of Conv groups, thereby further reducing the GroupNL model’s parameter count and FLOPs while maintaining performance.
+This paper also analyzes four types of NLFs: Sinusoidal, Monomial, Gaussian, and Laplace.
+It verifies across multiple datasets that the feature maps generated using the Sinusoidal function achieve the best performance.
+Finally, this paper conducts extensive experiments on the CIFAR-10, CIFAR-10-C, Icons-50, ImageNet-1K, and ImageNet-C datasets, and the experimental results demonstrate that our proposed GroupNL CNN models achieve strong robustness and resource-saving performance.
+For example, the accuracy of GroupNL ResNet-18 on Icons-50 dataset is about 2.86% higher than that of vanilla ResNet-18.
+On the ImageNet-C dataset, the accuracy of GroupNL-EfficientNet-ES is appropriately 1.1% higher than the vanilla EfficientNet-ES.
+In particular, with distributed data parallel training on ImageNet-1K, the GroupNL ResNet-101 achieves a 53% speedup over vanilla ResNet-101 while maintaining accuracy.
+We also deploy GroupNL models on resource-constrained devices (Raspberry Pi 4B), and on-device measurements show that the GroupNL models outperform in terms of inference speed, memory usage, power, and energy consumption.
 In summary, our main contributions are as follows.
-  • This paper proposes the GroupNL, which groups seed feature maps and generates the remaining feature maps using the NLF with different hyperparameters for each group.
-This improves the diversity of generated feature maps and the robustness of the CNN model without increasing its complexity.
-  • This paper introduces sparse technology to the ResNet-101 model, further reducing the number of learnable parameters and FLOPs of the model while maintaining model performance.
-In addition, it verifies that using the Sinusoidal function is the best to generate feature maps for GroupNL.
-  • Extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, and Icons-50, and ImageNet-1K datasets demonstrate that our proposed GroupNL achieves strong performance in terms of model robustness, number of learnable parameters, FLOPs, and training speedup.
+  • This paper proposes GroupNL, which groups seed feature maps and uses NLFs with fixed hyperparameters to generate diverse feature maps on demand for each Conv layer.
+This method improves the diversity of generated feature maps and enhances the CNN model’s robustness without increasing complexity.
+  • This paper proposes a sparse mode of GroupNL.
+By reasonably setting the relationship between the seed feature map ratio and the number of Conv groups, it further reduces the number of parameters and FLOPs of the GroupNL CNN model while ensuring performance.
+  • Extensive experiments on the CIFAR-10, CIFAR-10-C, and Icons-50, ImageNet-1K, and ImageNet-C datasets show that our proposed GroupNL achieves impressive performance in terms of model robustness, training/inference speedup.
 The remainder of the paper is organized as follows.
-Section 2 reviews the work on CNN and cloud-assisted CNN training and deployment.
+Section 2 reviews the work on cloud-assisted CNN training and deployment, and CNN.
 Section 3 describes the proposed GroupNL method in detail.
@@
 2 Related Work
-This section briefly introduces the Convolutional Neural Network (CNN) models and cloud-device CNN models training and deployment, which are inspiring our work.
-2.1 Convolutional Neural Network Models
-Because of the excellent feature extraction capabilities of the convolution operation, many popular CNN models based on convolution operations have been proposed.
-For example, Simonyan et al.
-[1] proposed the VGG model through stacked convolution operations with multiple model versions in different layers.
-He et al.
-[2] proposed the ResNet with residual branching, which solves the problem of degradation caused by deepening the convolutional layers.
-Ding et al.
-[4] introduced the re-parameterization technology to speed up the inference of the CNN model.
-To deploy CNN models in resource-constrained IoT devices, many lightweight CNN models have been proposed, such as MobileNet [13, 14, 15], ShuffleNet [16, 17], and GhostNet [18], etc.
-Some methods also consider the robustness of the CNN model, such as MonoCNN [5] and SineFM [19] for cloud-device collaborative deployment.
-TABLE I: Feature Maps Generated-based CNN Methods Training Evaluation with ResNet-101 in CIFAR-10 for 300 Epochs on NVIDIA RTX 2080Ti GPUs.
-Method Full Conv Nonlinear Func Cheap Linear Norm Top-1 Acc (%) Training Time
-Vanilla CNN 95.48 9.3 Hours
-MonoCNN 94.97 9.2 Hours
-GhostNet 95.04 12.7 Hours
-SineFM 95.06 15.4 Hours
-GroupNL (Ours) 95.51 8.3 Hours
-The most similar to our proposed GroupNL method are the GhostNet [18] and SineFM [19], which also generate feature maps via transformation.
-However, there are two significant differences as shown in Table I:
-(i) The GhostNet generates the feature maps with cheap linear transformation.
-The cheap linear is a variant of depthwise convolution (DWConv), in which the number of output channels is the multiple of the number of input channels, and the groups of convolution are equal to the number of input channels.
-So in this paper, we still refer to the cheap linear transformation as “DWConv”.
-Introducing trainable nodes, i.e., DWConv in the computational graph will result in slower training than vanilla CNN models.
-(ii) The SineFM mixes the nonlinear and cheap linear transformation for the feature maps generation, and also introduces batch normalization (Norm) for generated feature maps; SineFM has a better performance compared to GhostNet via nonlinear, nevertheless, it introduces trainable Norm and DWConv resulting in slower training.
-Compared to the above methods, our proposed GroupNL is designed with grouped nonlinear transformation to ensure the diversity of generated feature maps and better performance.
-At the same time, the generation in GroupNL only contains the copying and in-placing operators without extra trainable computational nodes, e.g., DWConv and Norm.
-2.2 Cloud-Device CNN Training and Deployment
+This section briefly introduces cloud-device training and deployment and the Convolutional Neural Network (CNN) models that inspire our work.
+2.1 Cloud-Device Training and Deployment
 Due to the limited resources of IoT devices, cloud-device collaborative training and deployment of CNN models have received widespread attention.
@@
 Cloud-Only.
-The cloud-only architecture is training and deploying the CNN models on the cloud server.
+The cloud-only architecture trains and deploys the CNN models on the cloud server.
 The user sends data and service requests to the cloud server, and after the cloud server processes the data, it sends the results to the users [12, 9, 19, 25].
@@
 They also process data on edge servers [14, 4, 1, 41, 5, 40] to reduce the amount of data uploaded to the cloud.
-However, the quality of services of cloud-only architecture is fully constrained by the network condition.
+However, the quality of services of cloud-only architecture is fully constrained by the network conditions.
 Cloud-Device Collaboration.
-The second architecture splits the CNN model into two parts, partial CNN layers are deployed on the cloud server and other parts on the IoT devices.
+The second architecture splits the CNN model into two parts: partial CNN layers are deployed on the cloud server, and other parts on the IoT devices.
 The cloud and IoT devices are collaborative training and inference of the split CNN model [20, 23, 22].
 For example, Kang et al.
-[30] proposed to divide the CNN into a head running on the device and a tail running on the cloud and decide the split point based on the load of the device and the conditions of the cloud and model.
+[20] proposed to divide the CNN into a head running on the device and a tail running on the cloud, and decide the split point based on the load of the device and the conditions of the cloud and model.
 Stefanos et al.
-[32] proposed a cloud-device collaborative inference method for the cloud and device with compression technologies to reduce the parameters exchange.
-However, cloud-device collaborative CNN methods have to search the optimal partitioning points labor-intensively for different CNN models.
+[22] proposed a cloud-device collaborative inference method for the cloud and device with compression technologies to reduce the parameter exchange.
+However, cloud-device collaborative methods have to search for the optimal partitioning points labor-intensively for different CNNs.
 Cloud-Assisted Training and Device Deployment.
-The third architecture is that the cloud server first trains the CNN model and then sends the trained CNN model to the IoT device for deployment [5, 6, 7, 33, 34].
+The third architecture is that the cloud server first trains the CNN model and then sends the trained CNN model to the IoT device for deployment [3, 45, 44, 38, 26, 43].
 We refer to it as the Cloud-Assisted Architecture.
@@
 The proposed GroupNL is based on the cloud-assisted training and IoT device deployment architecture.
-3 Design of the Proposed Approach
-3.1 Overview
-The proposed GroupNL is based on the cloud-assisted IoT device deployment and updating CNN model architecture.
-In this architecture, the CNN model is first trained in the cloud server, and then the trained CNN model is transferred to the IoT device for deployment and visual service.
-Two key factors motivate us to investigate a robust and low-resource CNN design method.
-(i) Typically, IoT devices such as smart surveillance and autonomous driving cameras work outdoors, and the CNN models deployed on the devices often process corrupted data affected by the natural environment, such as images collected on rainy, foggy, and snowy days.
-We refer to the performance of the on-device CNN model on corrupted data as model robustness.
-Most existing CNN models have low robustness when dealing with corrupted data.
-(ii) IoT devices are resource-constrained.
-To address the high resource consumption of vanilla CNN models, many alternatives to CNN have been proposed.
-Although alternatives to the vanilla CNN model reduce the number of model parameters and computation, they contain many additional operations (e.g., DWConv, Norm, etc.), which results in no significant acceleration of model training speed.
-To this end, we aim to design a novel, robust, and low-resource CNN model design method to improve the quality of service provided to users by cloud-assisted IoT devices in deploying and updating CNNs.
-Figure 2: An example of the GroupNL operation with $c_{\mathrm{in}}=8,c_{\mathrm{out}}=16$, and the number of seed features is $c_{\mathrm{seed}}=4$, the groups for splitting is $g=2$.
-In this case, the GroupNL can introduce $\gamma\cdot g=6$ different NLFs for the feature maps generation.
-Compared with the vanilla convolution operation needs 16 filters to compute all features, the GroupNL operation only requires 4 filters to compute the seed features and generates the remaining features by splitting, repeating, and diverse nonlinear transformations.
-3.2 Preliminaries of Nonlinear Transformation and Feature Maps Generated-based Methods in CNNs
-We illustrate our motivation by analyzing the most relevant works, nonlinear transformation method, i.e., MonoCNN [5], and feature maps generated-based methods, namely GhostNet [18] and SineFM [19], respectively.
-We also describe the operations of these CNN design methods.
-Given the input feature maps ${\bm{x}}$ and output feature maps ${\bm{y}}$ of this convolutional layer, where ${\bm{x}}\in\mathbb{R}^{h\times w\times c_{\mathrm{in}}}$ and ${\bm{y}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times w^{\!{}^{\prime}\!}\times c_{\mathrm{out}}}$.
-The $h$ and $w$ are the spatial dimensions of the input data/feature maps and $h^{{}^{\prime}}$ and $w^{{}^{\prime}}$ are the spatial dimensions of the output feature maps.
-$c_{\mathrm{in}}$ is the number of input channels and $c_{\mathrm{out}}$ is the number of output channels.
-For a detailed description of the nonlinear transformation and feature maps generated-based methods, we first introduce two concepts from these CNNs:
-(i) seed filters refer to a small number of convolutional filters in the convolution layer.
-(ii) seed feature maps refer to the feature maps generated by the seed filters and input via a convolution operation.
-We denote the seed filters are the ${\bm{W}}_{\mathrm{seed}}$, where ${\bm{W}}_{\mathrm{seed}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{seed}}}$ where $k$ is the spatial dimension of the filter assumed to be square.
+2.2 Convolutional Neural Network Models
+Because of the convolution operation’s excellent feature-extraction capabilities, many popular CNN models have been proposed.
+For example, Simonyan et al.
+[35] proposed the VGG model through stacked convolution operations with multiple model versions in different layers.
+He et al.
+[13] proposed the ResNet with residual branching, which solves the problem of degradation caused by deepening the convolutional layers.
+Ding et al.
+[6] introduced the re-parameterization technology to speed up the inference of the CNN model.
+To deploy CNN models in resource-constrained IoT devices, many lightweight CNN models have been proposed, such as MobileNet [16, 34, 17], EfficientNet [36, 8, 7, 37], ShuffleNet [46, 28], and GhostNet [11], etc.
+Some methods also consider the transmission cost and robustness of CNNs, such as SineFM [27] and MonoCNN [3] for cloud-device collaborative deployment.
+TABLE I: Comparison with the Existing Convolutions with Conv($n$,$m$,$k$,$g$), height is $h$, width is $w$, and $\xi=\mathrm{Gcd}(n,\frac{n}{r})$.
+Method Neural Network Modules #Ops #FLOPs
+Vanilla Conv($n$,$m$,kk,1) 1 $whmnk^{2}$
+Mono Conv($n$,$m$,kk,1) 1 $whmnk^{2}$
+[…]
+[…]
+[…]
+Depthwise Conv($n$,$n$,$k$,$n$) 1 $whn^{2}k^{2}\frac{1}{n}$
+[…]
+The most similar methods are Ghost-like Conv, i.e., GhostNet [11] and SineFM [27], which uses a seed Conv to get partial feature maps and leverage a transformation to generate the rest.
+  • The GhostNet generates the feature maps with cheap Conv Transformation [11].
+The “cheap Conv” is a variant of depthwise Conv [16], in which the number of output channels is a multiple of the number of input channels, and the groups of cheap Conv are equal to the common factor.
+  • The SineFM mixes nonlinear functions, Batch Normalization (BN), and cheap Conv for its feature maps generation.
+It performs better than GhostNet via nonlinearities; however, the introduced BN and cheap Conv result in slower computational speed.
+These Ghost-like Conv methods cannot achieve better acceleration because they use many Neural Network Modules, e.g., cheap Conv & BN.
+As shown in Table I, for a comparison, we propose GroupNL Conv to solve this issue without using cheap Conv and BN, which only uses a seed Conv with some lightweight Tensor Manipulation Operators like $\tt{torch.cat}$, $\tt{torch.split}$, and $\tt{torch.repeat}$, achieving a better accuracy and speed.
+3 The Proposed Method: GroupNL
+We introduce GroupNL, a low-resource and robust CNN design method for cloud-assisted IoT deployment, where models are trained in the cloud and deployed on resource-constrained devices.
+The GroupNL method addresses two key challenges: the demand for real-life resource-efficient training/inference, which prior lightweight alternatives have not fully met, and the need for robustness against real-world corruptions like adverse weather.
+We describe the proposed GroupNL in detail as follows: (i) preliminaries in Section 3.1, (ii) detailed design of GroupNL in Section 3.2, and (iii) training and inference acceleration analysis of GroupNL in Section 3.3.
+3.1 Preliminaries
+We illustrate our motivation by analyzing the most relevant works, the nonlinear transformation method MonoCNN [3], and Ghost-like Conv, including GhostNet [11] and SineFM [27], respectively.
+Given the input feature maps ${\bm{x}}\!\in\!\mathbb{R}^{h\times w\times c_{\mathrm{in}}}$ and output feature maps ${\bm{y}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times w^{\!{}^{\prime}\!}\times c_{\mathrm{out}}}$ of this Conv layer, where the $h$ and $w$ are the spatial dimensions of input feature maps, and $h^{{}^{\prime}}$ and $w^{{}^{\prime}}$ for output feature maps.
+$c_{\mathrm{in}}$, $c_{\mathrm{out}}$ are the number of input and output channels, respectively.
+For a detailed description, we introduce two concepts: (i) Seed filters refer to a small number of Conv filters, or referred to as seed Conv; (ii) Seed feature maps refer to feature maps generated by seed filters via seed Conv operation.
+We denote seed filters as ${\bm{W}}_{\mathrm{seed}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{seed}}}$, where $k$ is the kernel size.
 The number of seed filters is $c_{\mathrm{seed}}$, and the number of generated filters is $c_{\mathrm{gen}}$, where $c_{\mathrm{seed}}\!+\!c_{\mathrm{gen}}\!=\!c_{\mathrm{out}}$.
-We set the trainable computational operations belonging to $\tt{torch.nn.Module}$ in bold in equation, i.e., $\mathrm{\bf{DWConv}}$ and $\mathrm{\bf{Norm}}$.
-We also denote a reduction ratio $r$ representing the relationship between $c_{\mathrm{seed}}$ and $c_{\mathrm{out}}$:
-$\begin{array}[]{l}r={c_{\mathrm{out}}}/{c_{\mathrm{seed}}}\end{array}.$ (1)
+We denote an integer reduction ratio $r$ representing the relationship between $c_{\mathrm{seed}}$ and $c_{\mathrm{out}}$, i.e., $r={c_{\mathrm{out}}}/{c_{\mathrm{seed}}}$.
+We also set the trainable $\tt{torch.nn.Module}$ operations in bold in equation, e.g., $\mathrm{\bf{Conv}}$ and $\mathrm{\bf{BN}}$.
 MonoCNN [3].
-For a convolutional layer, some filters are designated as seed filters, and the remaining filters are generated based on the seed filters and the data-agnostic nonlinear transformation function (NLF).
-Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$, and a monomial NLF $\Psi_{\mathrm{mono}}(\cdot)$, we can obtain the remaining filters ${\bm{W}}_{\mathrm{gen}}$ of this layer as follows:
+For a Conv layer, some filters are designated as seed filters, and the remaining filters are generated based on seed filters and one data-agnostic nonlinear transformation function (NLF).
+Formally, given seed filters and a monomial NLF $\Psi_{\mathrm{mono}}(\cdot)$, the MonoConv obtains the remaining filters ${\bm{W}}_{\mathrm{gen}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{gen}}}$ as follows:
 $\begin{array}[]{l}{\bm{W}}_{\mathrm{gen}}=\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})\end{array},$ (1)
-where ${\bm{W}}_{\mathrm{gen}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{gen}}}$.
-Then, concatenating the seed filter ${\bm{W}}_{\mathrm{seed}}$ and the generated filters ${\bm{W}}_{\mathrm{gen}}$ to obtain the filters ${\bm{W}}$ of this layer:
+Then, concatenating the seed filter ${\bm{W}}_{\mathrm{seed}}$ and the generated filters ${\bm{W}}_{\mathrm{gen}}$ to obtain the filters ${\bm{W}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{out}}}$:
 $\begin{array}[]{l}{\bm{W}}=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};{\bm{W}}_{\mathrm{gen}}]\end{array},$ (2)
-where ${\bm{W}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{out}}}$.
-Finally, perform the vanilla convolution operation on the input data/feature maps to obtain the output ${\bm{y}}$ of this layer:
-${\bm{y}}$ $={\bm{W}}*{\bm{x}}$ (4)
-$=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})]*{\bm{x}}\ ,$
-where * is the convolution operation.
-The MonoCNN improves the robustness of the CNN model via data-agnostic NLFs for filters but still with a complete convolution operation.
-GhostNet [18]. The GhostNet introduces partial convolution filters ${\bm{W}}_{\mathrm{seed}}$ and generated the seed feature maps as follows:
-$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}={\bm{W}}_{\mathrm{seed}}*{\bm{x}}\end{array},$ (5)
+Finally, perform a vanilla Conv operation on input feature maps to obtain the output $\bm{y}$} of MonoConv:
+${\bm{y}}={\bm{W}}*{\bm{x}}=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})]*{\bm{x}}\ ,$ (3)
+where * is the Conv operation. The MonoCNN improves the robustness of the CNN model via data-agnostic NLFs for filters, but still with a complete Conv operation.
+GhostNet [11]. The Ghost Conv introduces $k\times k$ seed Conv to generate seed feature maps:
+$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}={\bm{W}}_{\mathrm{seed}}*{\bm{x}}=\mathop{\mathrm{\bf{Conv}}}\limits_{\ k\times k}({\bm{x}})\end{array},$ (4)
 where ${\bm{y}}_{\mathrm{seed}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times c_{\mathrm{seed}}}$.
-For the remaining feature maps, GhostNet introduces additional cheap linear transformation, i.e., the $d\times d$ depthwise convolution filters to process the ${\bm{y}}_{\mathrm{seed}}$ for generating features:
-${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (6)
-$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{d\times d}({\bm{y}}_{\mathrm{seed}})]\ .$
-However, the cheap linear operation (i.e., DWConv op.) introduced by GhostNet causes a slow computation in training.
-At the same time, the parameters of DWConv are also trainable, so the robustness of GhostNet is completely restricted to the training data.
+For the remaining feature maps, GhostNet introduces additional cheap linear transformation, i.e., the $d\!\times\!d$ cheap Conv to process the ${\bm{y}}_{\mathrm{seed}}$ for generating features:
+${\bm{y}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{Conv}}}\limits_{d\times d}({\bm{y}}_{\mathrm{seed}})]\ .$ (5)
+However, the cheap Conv introduced by GhostNet slows down training and inference.
+Also, the parameters of cheap Conv are trainable, so GhostNet’s robustness is completely restricted to the training data.
 SineFM [27].
-Similar to the GhostNet, some filters of SineFM are designated as seed filters for a convolution layer.
-First, seed feature maps equal to the number of seed filters are generated based on the seed filters and input data.
-Then, feature maps with $t$-times the number of $c_{\mathrm{seed}}$ are generated based on the seed feature maps and the NLFs.
-Finally, using $1\times 1$ depthwise convolution for channel alignment.
-Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer by the Eq.
-5.
-Then, $t$-data-agnostic NLFs $\bm{{\Psi}}_{\mathrm{sinefm}}={\{\Psi_{i}\|i=1,\dots,t\}}$ process the seed feature maps to obtain more features:
-$\begin{array}[]{l}{\bm{y}}_{\mathrm{exp}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{exp}_{1}};\cdots;{\bm{y}}_{\mathrm{exp}_{t}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[{\bm{y}}_{\mathrm{exp}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{exp}_{i}}=\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))\ ,\end{array}$ (7)
-where ${\bm{y}}_{\mathrm{exp}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!c_{\mathrm{seed}}}$, ${\bm{y}}_{\mathrm{exp}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\times\!t\cdot c_{\mathrm{seed}}}$, and $i\!=\!1,\!\cdots,\!t$.
-The $t$ is the times of expanding seed feature maps.
-The generated feature maps expanding by $t$ times of seed feature maps will usually exceed the number of feature maps required for this layer.
-Each expanding seed feature maps ${\bm{y}}_{\mathrm{exp}_{i}}$ also process a specific $\mathrm{\bf{Norm}}_{i}$.
-All expanding seed feature maps are concatenated and then process the $1\times 1$ depthwise convolution filters to align the dimensions and obtain the generated feature maps:
-$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}({\bm{y}}_{\mathrm{exp}})\ .\end{array}$ (8)
+For a SineFM Conv, similar to the Ghost Conv, given seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps $\bm{x}$}, we can obtain seed feature maps ${\bm{y}}_{\mathrm{seed}}$ by the Eq.
+4.
+Then, $t$-NLFs $\bm{{\Psi}}_{\mathrm{sinefm}}\!=\!{\{\Psi_{i}\|i\!=\!1,\!\dots\!,t\}}$ process the seed feature maps to obtain expanded feature maps ${\bm{y}}_{\mathrm{exp}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!t\cdot c_{\mathrm{seed}}}$ with more features:
+$\begin{array}[]{l}{\bm{y}}_{\mathrm{exp}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{exp}_{1}};\cdots;{\bm{y}}_{\mathrm{exp}_{t}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[{\bm{y}}_{\mathrm{exp}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{exp}_{i}}=\mathrm{\bf{BN}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))\ ,\end{array}$ (6)
+where ${\bm{y}}_{\mathrm{exp}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!c_{\mathrm{seed}}}$, and $i\!=\!1,\!\cdots\!,\!t$.
+The $t$ is the time of expanding seed feature maps.
+The expanded feature maps will exceed the number of feature maps required for this layer.
+Each expanded seed feature maps ${\bm{y}}_{\mathrm{exp}_{i}}$ also process a specific $\mathrm{\bf{BN}}_{i}$.
+Then, expanded feature maps are processed by $1\!\times\!1$ cheap Conv to align the $c_{\mathrm{gen}}$ dimensions:
+$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathop{\mathrm{\bf{Conv}}}\limits_{1\times 1}({\bm{y}}_{\mathrm{exp}})\ .\end{array}$ (7)
 Finally, the generated feature maps are concatenated with the seed feature maps to form the output of the SineFM layer as follows:
 ${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (8)
-$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}(\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))])]\ .$
+$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{Conv}}}\limits_{1\times 1}(\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[\mathrm{\bf{BN}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))])]\ .$
+Figure 2: The example of GroupNL Conv with $c_{\mathrm{in}}\!=\!8,c_{\mathrm{out}}\!=\!16$, and the number of seed features is $c_{\mathrm{seed}}\!=\!4$, the groups for splitting is g=2g\!=\!2.
+In this case, the GroupNL can introduce $\gamma\cdot g\!=\!6$ different NLFs for the feature maps generation.
+Compared with the vanilla Conv, which needs 16 filters to compute all features, the GroupNL Conv only requires 4 filters to compute the seed features and generates the remaining features by nonlinearities and tensor manipulation operators.
 3.2 Design of GroupNL
-3.3.1 Grouped Nonlinear Transformation Algorithm
-The aforementioned GhostNet and SineFM methods introduce too much computation in feature maps generation by DWConv and Norm operations which result in more computational graph nodes and slow down training.
-At the same time, the data-agnostic nonlinear transformation of SineFM is only $t$-different types.
-Therefore, resource efficiency and robustness are needed to be further improved.
-The key challenge is how to increase the diversity of data-agnostic nonlinear transformations to improve the robustness of the model without increasing the resource computation of running the model.
-To this end, we propose the Grouped NonLinear transformation feature map generation (GroupNL) method, which is a robust and low-resource CNN model design method.
-Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input data/feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer as same as the Eq.
-5.
-Besides, we designate the seed filters ${\bm{W}}_{\mathrm{seed}}$ as the standard convolution in basic structures like ResNet-18/-34 and VGG11, and sparse in bottleneck structures, e.g., ResNet-101.
-The sparse filter means that the number of groups of the convolution operation of seed filters is the Greatest Common Divisor (GCD) of $c_{\mathrm{in}}$ and $c_{\mathrm{seed}}$, i.e., $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$.
-Introducing the sparse seed convolution in the larger bottleneck structure can further improve the resource efficiency of the larger CNN models.
-Utilizing the standard seed convolution in basic structure is beneficial for maintaining the performance of the smaller CNN models.
-Splitting and Grouped Algorithm. We split the seed feature maps into $g$ groups in $c_{\mathrm{seed}}$ dimension instead of expanding, as follows:
+3.2.1 Grouped Nonlinear Transformation
+The aforementioned GhostNet and SineFM methods introduce too much computation in feature maps generation by cheap Conv and BN operations, which result in slow training/inference speed.
+At the same time, NLFs in SineFM are only $t$-different types.
+Therefore, resource efficiency and robustness need to be further improved.
+The key challenge is how to increase the diversity of data-agnostic NLFs to improve the robustness of the model without increasing computations.
+To this end, we propose the Grouped NonLinear transformation (GroupNL) method.
+Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps $\bm{x}$}, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ with the Eq.
+4.
+Splitting and Grouped Algorithm. We split seed feature maps into $g$ groups in $c_{\mathrm{seed}}$ dimension instead of expanding:
 $\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}_{1}};\cdots;{\bm{y}}_{\mathrm{seed}_{g}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[{\bm{y}}_{\mathrm{seed}_{i}}]\end{array},$ (9)
-where $g$ is a positive integer, and the modulo result is $\mathrm{Mod}(c_{\mathrm{seed}},g)\!=\!0$, and ${\bm{y}}_{\mathrm{seed}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\frac{c_{\mathrm{seed}}}{g}},i=1,\cdots,g$.
+where $g$ is a positive integer, where its modulo result is $\mathrm{Mod}(c_{\mathrm{seed}},g)\!=\!0$, and ${\bm{y}}_{\mathrm{seed}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\frac{c_{\mathrm{seed}}}{g}},i=1,\!\cdots\!,\!g$.
 For the remaining dimension $c_{\mathrm{gen}}\!=\!c_{\mathrm{out}}\!-\!c_{\mathrm{seed}}$, we want to represent it by $c_{\mathrm{seed}}$ without expanding the dimension.
-We denote a multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=(r\cdot c_{\mathrm{out}}/c_{\mathrm{in}}-1)$, so each group of seed feature maps should be copied $\gamma$ times for filling the remaining dimension.
+We denote a multiplier $\gamma\!=\!\frac{c_{\mathrm{gen}}}{c_{\mathrm{seed}}}\!=\!r\cdot\frac{c_{\mathrm{out}}}{c_{\mathrm{in}}}-1$, so each group of seed feature maps should be copied $\gamma$ times for filling the remaining dimension.
 Thus, the total number of copied groups of feature maps is $\gamma\cdot g$.
@@
 By splitting and copying, we have $\gamma\cdot g$ groups of identical seed feature maps, and no additional dimensions are introduced.
-We can introduce $\gamma\cdot g$ data-agnostic different nonlinear transformations, i.e., diverse NLFs $\bm{{\Psi}}_{\mathrm{groupnl}}={\{\Psi_{i,j}\|i=1,\dots,g;j=1,\dots,\gamma\}}$ to these identical groups.
-Subsequently, for each copied group of seed feature maps, NLF with different hyperparameter settings is used to generate the corresponding feature maps, as follows:
+We can introduce $\gamma\cdot g$ data-agnostic different nonlinear transformations, i.e., diverse NLFs $\bm{{\Psi}}_{\mathrm{gnl}}\!=\!{\{\Psi_{i,j}\|i\!=\!1,\dots,g;j\!=\!1,\dots,\gamma\}}$ to these identical groups.
+Subsequently, for each copied group of seed feature maps, NLFs with different hyperparameter settings are used to generate corresponding feature maps:
 $\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{gen}_{1}};\cdots;{\bm{y}}_{\mathrm{gen}_{g}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[{\bm{y}}_{\mathrm{gen}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{gen}_{i}}=\mathop{\mathrm{Concat}}\limits_{j=1,\cdots,\gamma}[\Psi_{i,j}({\bm{y}}_{\mathrm{seed}_{i}})]\ ,\end{array}$ (10)
@@
 2, we can describe the operations of grouped diverse nonlinear transformation in detail with an example of $c_{\mathrm{in}}\!=\!8,c_{\mathrm{out}}\!=\!16,c_{\mathrm{seed}}\!=\!4,g\!=\!2$.
-In this case, the number of generated features $c_{\mathrm{gen}}=16-4=12$, the multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=12/4=3$, and introducing $|\bm{{\Psi}}_{\mathrm{groupnl}}|=\gamma\cdot g=6$ different NLFs for feature maps generation.
-Finally, generated feature maps are concatenated with the seed feature maps to produce the output of the GroupNL layer, as follows:
+In this case, the number of generated features $c_{\mathrm{gen}}\!=\!16\!-\!4\!=\!12$, the multiplier $\gamma\!=\!c_{\mathrm{gen}}/c_{\mathrm{seed}}\!=\!12/4\!=\!3$, and introducing $|\bm{{\Psi}}_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!6$ different NLFs for feature maps generation.
+GroupNL Conv. Finally, generated feature maps are concatenated with the seed feature maps to produce the output of the standard GroupNL Conv layer, as follows:
 ${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (11)
 $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,g}[\mathop{\mathrm{Concat}}\limits_{j=1,\cdots,\gamma}[\Psi_{i,j}({\bm{y}}_{\mathrm{seed}_{i}})]]]\ .$
-The PyTorch-like pseudocode of the GroupNL-based convolutional layer is demonstrated in Algorithm 1.
-As described in Eq.
-12 and Algorithm 1, the generating procedure of GroupNL is all in-placing (e.g., splitting $\tt{torch.split}$, concatenation $\tt{torch.cat}$) and copying (e.g., $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$) operators, and data-agnostic and random nonlinear transformation operations, without any extra $\tt{torch.nn.Module}$ (i.e., DWConv and Norm) compared with the GhostNet in Eq.
+Sparse GroupNL Conv.
+In addition to the standard GroupNL Conv, we introduce a more lightweight module, termed Sparse GroupNL Conv.
+In this design, the seed filters ${\bm{W}}_{\mathrm{seed}}$ are grouped for thinning, where the number of groups $\xi$ in seed Conv is set to the Greatest Common Divisor (GCD) between $c_{\mathrm{in}}$ and $c_{\mathrm{seed}}$, i.e., $\xi\!=\!\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$.
+The number of parameters, FLOPs, and energy consumption in Sparse GroupNL Conv are close to or slightly lower than depthwise Conv, while offering higher computational efficiency, as profiling in Section 4.4.3.
+Therefore, Sparse GroupNL Conv can be used to replace depthwise Conv in lightweight models, e.g., MobileNet, EfficientNet, to improve the robustness.
+Meanwhile, we observe that Sparse GroupNL Conv preserves performance in large-scale CNNs with bottleneck structures (e.g., ResNet-101), whereas the standard GroupNL Conv is still required for compact models based on basic blocks (e.g., ResNet-18/34).
+The PyTorch-like pseudocode of the GroupNL Conv layer is demonstrated in Algorithm 1.
+As shown, its generating procedures are all tensor manipulation operators, e.g., splitting $\tt{torch.split}$, concatenation $\tt{torch.cat}$, copying $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, and data-agnostic and random nonlinear transformation operations, without any extra $\tt{torch.nn.Module}$ (Conv and BN) compared with GhostNet in Eq.
 5 and SineFM in Eq.
-9.
+8.
 # func: nonlinear transformation function
@@
 # C_i, C_s, C_o: No. of input/seed/output features
-class GroupNLConv2d(nn.Conv2d):
+class GroupNLConv2d($n$.Conv2d):
 def __init__(self, C_in, C_out, C_s, g, func, sparse, **kwargs):
@@
 C_g = C_s * (ceil(C_o/C_s) - 1)
-# seed filters for generating feasures
-self.conv = nn.Conv2d(C_i, C_s, \
+# seed filters for generating features
+self.conv = $n$.Conv2d(C_i, C_s, \
 groups=math.gcd(C_i, C_s) \
@@
 # diverse random hyperparameters of func
-self.hy=nn.Parameter(torch.rand(gamma*g))
+self.hy=$n$.Parameter(torch.rand(gamma*g))
 # ensure non-learnable hyperparameters
@@
 return torch.cat([y_s, y_g])
-Algorithm 1 GroupNL-based Convolutional Layer: PyTorch-like Pseudocode
-3.3.2 Diversities of Data-Agnostic Nonlinear Transformation Function and Feature Maps Generation
-The number of NLFs in GroupNL is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g$, which is usually larger than the fixed number of $t$ NLFs in SineFM $|\Psi_{\mathrm{sinefm}}|=t$, which act on all seed feature maps for expanding, limiting the diversity of features.
-For the diversities of data-agnostic NLFs, assuming there are $64$ seed feature maps, and the number of output channels is $c_{\mathrm{out}}=256$, we can use different NLFs to generate the remaining $192$ feature maps.
-Assuming $t=5$, SineFM needs to generate $t\cdot c_{\mathrm{seed}}=5\cdot 64=320$ expanding features and utilizes a DWConv to align them into $192$ features.
-In this case, the number of NLFs is $5$.
+Algorithm 1 GroupNL Conv: PyTorch Pseudocode
+3.2.2 Diversities of Nonlinearities and Feature Maps
+The number of NLFs in GroupNL is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g$, which is usually larger than the fixed number of $t$ NLFs in SineFM $|\Psi_{\mathrm{sinefm}}|\!=\!t$, which act on all seed feature maps to expand or limit feature diversity.
+Diversities of Data-agnostic NLFs.
+Assuming there are $64$ seed feature maps, and the number of output channels is $c_{\mathrm{out}}\!=\!256$, we can use different NLFs to generate the remaining $192$ feature maps.
+For SineFM, assuming t=5t\!=\!5, it needs to generate $t\cdot c_{\mathrm{seed}}\!=\!5\cdot 64\!=\!320$ expanding features and utilizes a cheap Conv to align them into $192$ features.
+So, the number of NLFs is $5$.
 For GroupNL, the $\gamma\!=\!c_{\mathrm{gen}}/c_{\mathrm{seed}}\!=\!192/64\!=\!3$.
-For $g=2$, the number of different NLFs is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 2=6$; for $g=4$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 4=12$; for $g=8$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 8=24$; and for the fully grouped $g=64$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 64=192\gg|\Psi_{\mathrm{sinefm}}|=5$.
+For $g=2$, the number of different NLFs is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 2\!=\!6$; for g=4g\!=\!4 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 4\!=\!12$; for g=8g\!=\!8 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 8\!=\!24$; and for the fully grouped g=64g\!=\!64 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 64\!=\!192\!\gg\!|\Psi_{\mathrm{sinefm}}|\!=\!5$.
 Compared with the SineFM, the diversity of NLFs is significantly improved without increasing the number of feature maps in GroupNL.
-For the diversities of generated feature maps, we can visualize a given input image processing with 24-different NLFs as an example, as shown in Fig.
+Diversities of Generated Feature Maps.
+We can visualize a given input image processing with 24-different NLFs as an example, as shown in Fig.
 3.
 The richness of the generated feature maps is greatly improved while maintaining the similarity between different features.
-Figure 3: The visualization of feature maps generated based on the diverse NLFs with 24-different hyperparameter settings.
+Figure 3: The visualization of feature maps generated based on the diverse NLFs with 24 different hyperparameter settings.
 The diversity of the generated feature maps is greatly improved while maintaining the similarity.
-3.3.3 Selection of NLFs, Hyperparameters and Coefficients
+3.3 Training & Inference Acceleration Analysis
+Cloud Training Acceleration.
+For cloud-assisted CNN models, the multi-GPU training on cloud servers is mainly based on data parallelism.
+In practical multi-GPU training, such as Data Parallelism (DP) [32] and Distributed Data Parallelism (DDP) [24], the bandwidth of inter-GPU memory data transfers is an important constraint on the training speed compared to the powerful and fast computation kernels.
+So, the number of gradients can be further optimized by reducing inter-GPU scatter and by optimizing different data parallelism levels.
+Assuming each trainable parameter in a CNN model maintains a gradient, the number of gradients $\mathcal{G}$ corresponds to the number of trainable parameters $\mathcal{P}$.
+For $N$-GPUs DP, one iteration data communication in the main GPU is $(N\!-\!1)\mathcal{G}$; for the rest N1N-1 GPUs are $(N\!-\!1)\mathcal{G}$, where each GPU sends $\mathcal{G}$.
+For $N$-GPUs DDP, one iteration data communication in a single GPU is $\frac{2(N-1)}{N}\mathcal{G}$.
+Similar to lightweight models, GroupNL essentially achieves training acceleration by reducing the $\mathcal{P}$ (and thus the $\mathcal{G}$).
+Based on this premise, we analyze the theoretical gradient exchange in multi-GPUs memory of vanilla CNN, GhostNet, SineFM, and GroupNL in training.
+Ignoring the bias, for a vanilla Conv layer, the number of parameters and gradients can be described as follows:
+$\begin{array}[]{l}{\mathcal{P}_{\mathrm{conv}}=\mathcal{G}}_{\mathrm{conv}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\ .\end{array}$ (12)
+For a Ghost Conv layer, the number of parameters and gradients from the seed Conv and $d\!\times\!d$ cheap Conv are:
+${\mathcal{P}_{\mathrm{ghost}}=\mathcal{G}}_{\mathrm{ghost}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+c_{\mathrm{seed}}c_{\mathrm{gen}}d^{2}\frac{1}{c_{\mathrm{seed}}}$ (13)
+$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r}+c_{\mathrm{out}}\frac{r\!-\!1}{r}d^{2}\ ,$
+and the number of parameters and gradients for the SineFM from the seed Conv, $1\!\times\!1$ cheap Conv, and BN, are:
+${\mathcal{P}_{\mathrm{sinefm}}=\mathcal{G}}_{\mathrm{sinefm}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+tc_{\mathrm{seed}}c_{\mathrm{gen}}\frac{1}{c_{\mathrm{seed}}}+2tc_{\mathrm{seed}}$ (14)
+$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r}+c_{\mathrm{out}}\frac{t(r\!+\!1)}{r}\ .$
+The proposed GroupNL Conv layer only contains the learnable standard seed filters and learnable sparse seed filters, assuming $\xi=\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$, and the number of parameters and gradients are shown in Eq.
+15:
+${\mathcal{P}_{\mathrm{gnl}}=\mathcal{G}}_{\mathrm{gnl}}$ $={c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}}={c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r}}\ ,$ (15)
+$\mathcal{P}_{\mathrm{gnl}}^{\mathrm{sp}}=\mathcal{G}_{\mathrm{gnl}}^{\mathrm{sp}}$ $={c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}}\frac{1}{\xi}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r\xi}\ .$
+Noticeably, the relation is $\mathcal{G}_{\mathrm{gnl}}^{\mathrm{sp}}\!\leq\!\mathcal{G}_{\mathrm{gnl}}\!=\!c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r}\!<\!\mathrm{min}\{{\mathcal{G}}_{\mathrm{conv}},{\mathcal{G}}_{\mathrm{ghost}},{\mathcal{G}}_{\mathrm{sinefm}}\}$, i.e., the number of gradients for optimization in GroupNL is minimal compared to the rest CNN design methods of ${\mathcal{G}}_{\mathrm{conv}}$, ${\mathcal{G}}_{\mathrm{ghost}}$, and ${\mathcal{G}}_{\mathrm{sinefm}}$.
+The sparse GroupNL also significantly reduces the gradient exchange by a factor of ${(r\xi)}^{-1}$ compared to the ${\mathcal{G}}_{\mathrm{conv}}$.
+Notably, it seems that Ghost Conv and SineFM also have a small number of gradients with ${{\mathcal{G}}_{\mathrm{ghost}}\!<\!\mathcal{G}}_{\mathrm{conv}}$ and ${\mathcal{G}}_{\mathrm{sinefm}}\!<\!{\mathcal{G}}_{\mathrm{conv}}$, can also accelerate training in theory.
+However, the evaluation in Table VIII shows that the GhostNet and SineFM are even slower than the vanilla CNN.
+For this phenomenon, we consider this to be due to the introduction of more computational nodes involved in the backward updating, such as cheap Conv and BN, which leads to slower training instead.
+The GroupNL, on the other hand, does not introduce new computational nodes through tensor manipulation operations such as $\tt{torch.cat}$, $\tt{torch.split}$, $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, thus also ensures acceleration of training.
+On-Device Inference Efficiency.
+For a fixed feature map size $wh$, the computational cost in FLOPs is proportional to the number of parameters, resulting in $\mathcal{F}_{\mathrm{gnl}}^{\mathrm{sp}}\!\leq\!\mathcal{F}_{\mathrm{gnl}}\!=\!c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\frac{1}{r}\!<\!\mathrm{min}\{{\mathcal{F}}_{\mathrm{conv}},{\mathcal{F}}_{\mathrm{ghost}},{\mathcal{F}}_{\mathrm{sinefm}}\}$, confirming that GroupNL Conv requires significantly fewer computations than other Convs.
+Furthermore, GroupNL employs the fewest $\tt{torch.nn.Module}$, contributing to its theoretically superior inference efficiency on devices.
+The module-level profiling in Section 4.4.3 also demonstrates the practical inference efficiency.
+3.4 Discussion
+As the aforementioned analysis, compared with the GhostNet [11] and SineFM [27], the proposed GroupNL-based CNN design method has two highlights:
+(i) More diverse feature maps.
+Given the seed filters and the number of output channels of the layer, by grouping the seed filters, more different hyperparameter settings of data-agnostic NLFs can be used to generate diverse features.
+(ii) More efficient training.
+GroupNL generates the number of feature maps without extra computational nodes like cheap Conv and BN, effectively reducing the number of gradients exchanging in the multi-GPUs training.
+4 Experiments
+This section introduces our experimental setup, including the datasets, baselines, and implementation details.
+Then, we empirically compare accuracy, network traffic, FLOPs, and training speed on multiple benchmarks.
+4.1 Experimental Setup
+4.1.1 Datasets
+We evaluate the effectiveness of the proposed method using five widely used datasets.
+CIFAR-10 and ImageNet-1K are standard datasets.
+CIFAR-10-C, Icons-50, and ImageNet-C are used as corrupted datasets.
+CIFAR-10 [21] is a multi-class natural object dataset for image classification.
+It consists of 50,000 training images and 10,000 test images in 10 categories, each with a resolution of $32\!\times\!32$ pixels.
+CIFAR-10-C [15] is a test dataset after using synthetic common perturbations and noise corruptions on the CIFAR-10 test set.
+It consists of 10,000 test images of 19 types of damage in 4 categories.
+The resolution of each image is $32\!\times\!32$ pixels.
+Icons-50 [15] consists of 10,000 images in 50 categories collected from different companies.
+We set the resolution of all images to $32\!\times\!32$ pixels.
+For training, the data of one company is retained as test data, and the data of the rest of the companies is used as training data.
+Please refer to [3] for more explanation of the CIFAR-10-C and Icons-50 datasets.
+ImageNet-1K [2] is a large-scale classification dataset to verify computer vision performance.
+ImageNet-1K is a subset of the ILSVRC ImageNet, which contains 1.28 billion training images and 50K validation images from 1K different classes.
+The resolution of each image is $224\!\times\!224$ pixels.
+ImageNet-C [15] is a large-scale dataset designed to evaluate the robustness of CNNs against common corruptions.
+The ImageNet-C includes 19 types of corruptions at five severity levels (as shown in Figure 4) in 5 categories: Noise, Blur, Weather, Digital, and Extra, simulating real-world challenges, e.g., environmental conditions, based on ImageNet-1K validation data.
+The 19 types include: gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness, contrast, elastic transformation, pixelate, JPEG compression, speckle noise, gaussian blur, spatter, and saturate.
+The resolution of the corrupted image is $224\!\times\!224$.
+Figure 4: Visualization examples of different severity levels of corruptions on ImageNet-C.
+4.1.2 Baseline CNNs
+We compare the following robust and feature maps generation CNNs discussed in Section 3.1: MonoCNN [3], GhostNet [11], and SineFM [27], and the above baseline methods are based on the standard CNN models: ResNet-18/-34/-101 [13] and VGG11 [35].
+For evaluations of large-scale datasets, i.e., ImageNet-1K and ImageNet-C, we choose representative lightweight CNNs widely used in mobile edge AI, including MobileNet-V2 [34], MobileNet-V3 [17], EfficientNet-Lite0 [8], EfficientNet-EdgeTPU-Small [7] (EfficientNet-ES), EfficientNet-V2-Small [37].
+To make the effect of training acceleration more clear, we evaluate the training speed of a larger model, i.e., ResNet-101.
+For on-device resource measurements, we deploy vanilla and GroupNL-based MobileNetV3, EfficientNet-ES, ResNet-50, and ResNet-101.
+4.1.3 Implementation Details
+The implementation details include training platforms & libraries and experimental settings.
+Training Platforms and Libraries: The experiments are performed on the cloud servers with NVIDIA RTX 2080Ti 11GB and RTX 4090 GPUs 24GB.
+We implement the models in PyTorch 1.12 for the NVIDIA platforms with CUDA 11.1.
+Hyperparameters Settings of Baseline CNNs: For the hyperparameters setting, we follow the suggestion from the original papers [3, 11, 27].
+The specific description of model hyperparameters is shown in Table II.
+  • MonoCNN [3] uses the monomial function as a filter generating function and samples a random and continuous exponent $\eta$ between $[1,7]$, and the exp_factor $\bm{e}$ is $\{2,4,8,16\}$;
+  • GhostNet [11] uses a d=3d\!=\!3 grouped convolution as the cheap operation for generating feature maps.
+  • SineFM [27] uses a sinusoidal function, the ranges of period $\omega$ and shift $\phi$ are between $[1,2]$ and $[1,5]$, respectively. The number of NLFs in SineFM is t=5t\!=\!5.
+We also set up the same scale of seed feature map numbers in GhostNet, SineFM, and GroupNL with r=2r\!=\!2.
+Experimental Settings: For the settings of training in CIFAR-10/Icons-50, the learning rate is 0.1, and the batch size is 128, with the SGD optimizer and cosine annealing scheduler.
+For training ResNet architectures in ImageNet-1K, the epochs is 300, the learning rate is 0.001, and the batch size is 1024, with the Adam optimizer and cosine annealing scheduler.
+For training lightweight CNNs in ImageNet-1K, the learning rate is 0.064, and the batch size is 1024, with the RMSprop optimizer and step scheduler.
+For MobileNet-V2, EfficientNet-Lite0, EfficientNet-ES, and EfficientNet-V2-S, the epoch is 450; for MobileNet-V3, the epoch is 600.
+In addition, the training evaluation is compared for the larger model, i.e., ResNet-101, in different parallel technologies in PyTorch [32], including DP [32], DDP [24], and Automatic Mixed Precision (AMP) [29].
+We also follow the linear scaling rule [10] to scale the number of parallelism in training, which is scaling the learning rate by the batch size.
+Edge Device Settings: For hardware, we deploy a cloud-device inference system demo based on our cloud server and a resource-constrained Raspberry Pi 4B with a 4-Core Broadcom BCM2711 ARM Cortex-A721.50GHz CPU and 8GB RAM, and use the USB-C power meter, POWER-Z KM003C, to record the power and energy changes on Raspberry Pi 4B during inference, as shown in Fig.
+5.
+For on-device libraries, we choose two inference backends, i.e., PyTorch-CPU [32] and ONNX [31]+ONNX Runtime [30], for deployment on Raspberry Pi. The versions are ONNX 1.14.1 and ONNX Runtime 1.15.1.
+TABLE II: Hyperparameter setting of CNNs methods.
+Method Reduction, Coefficient & Range
+MonoCNN $\bm{e}=\{2,4,8,16\},\eta\in[1,7]$
+GhostNet $r=2,d=3$
+SineFM $r=2,t=5,\omega\in[1,2],\phi\in[1,5]$
+GroupNL $r=2,g=4,\omega\in[1,2],\phi\in[1,5]$
+Figure 5: Hardware: a Resource-constrained Raspberry Pi 4B for Deployment and a Power Meter for Energy Evaluation.
+4.2 Experimental Results
+TABLE III: Performance Comparison on CIFAR-10.
+Method #Params (M) #FLOPs (M) Top-1 Acc (%)
+VGG11 Vanilla CNN 9.23 152.92 91.89
+MonoCNN 0.66 162.18 88.84
+GhostNet 4.64 77.22 92.65
+SineFM 4.64 77.68 91.27
+GroupNL 4.62 76.62 91.62
+ResNet-18 Vanilla CNN 11.17 556.04 95.04
+MonoCNN 2.58 565.51 92.05
+GhostNet 5.61 281.13 93.36
+SineFM 5.57 280.81 94.36
+GroupNL 5.60 279.49 95.12
+ResNet-34 Vanilla CNN 21.28 1160.43 95.24
+MonoCNN 3.66 1180.05 92.26
+GhostNet 10.68 584.97 93.90
+SineFM 10.65 586.28 94.77
+GroupNL 10.65 582.09 95.45
+ResNet-101 Vanilla CNN 42.51 2515.09 95.48
+MonoCNN 16.33 2544.66 94.97
+GhostNet 30.51 1838.52 95.04
+SineFM 29.48 1796.49 95.06
+GroupNL 18.46 1158.68 95.51
+TABLE IV: Performance Comparison on ImageNet-1K.
+Model #Params (M) #FLOPs (M) Accuracy (%)
+Top-1 Top-5
+MobileNet-V2 3.50 307.45 70.0 89.4
+MobileNet-V2-GroupNL 3.51 308.60 69.8 89.3
+MobileNet-V3 5.48 221.59 73.4 91.6
+MobileNet-V3-GroupNL 5.48 222.28 72.6 90.9
+EfficientNet-Lite0 4.65 391.93 72.8 91.4
+EfficientNet-Lite0-GroupNL 4.65 393.08 72.5 91.1
+EfficientNet-ES 5.44 1797.46 75.9 93.2
+EfficientNet-ES-GroupNL 5.44 1798.28 75.9 93.3
+EfficientNet-V2-S 21.46 2865.40 83.0 96.2
+EfficientNet-V2-S-GroupNL 21.46 2867.05 83.1 96.3
+This section evaluates the robustness of the GroupNL-based CNN on three types of datasets, including standard datasets (i.e., CIFAR-10 and ImageNet-1K), the corrupted dataset (i.e., CIFAR-10-C and ImageNet-C), and data under different styles (i.e., Icons50).
+4.2.1 Evaluation on Standard Datasets
+Table III and Table IV show the comparison of image classification results of different CNN models on CIFAR-10 and ImageNet-1K datasets under different methods.
+We observe that GroupNL-based CNNs generally outperform the accuracy of alternatives to standard CNN models.
+As shown, when using ResNet-34 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.68% higher than SineFM’s, 1.55% higher than GhostNet’s, and 3.19% higher than MonoCNN’s.
+When using VGG11 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.35% higher than SineFM’s.
+The main reason is that GroupNL generates diversified feature maps using the NLFs with different hyperparameter configurations.
+In contrast, SineFM generates feature maps using a single NLF and is unable to generate diverse feature maps.
+GhostNet first generates some feature maps using standard Conv operations and then uses cheap linear operations to generate other feature maps based on these generated feature maps, which limits the diversity of the feature maps.
+Benefiting from its ability to generate diverse feature maps, GroupNL achieves the highest accuracy among many alternatives.
+To maintain the number of model parameters and FLOPs, the lightweight CNN (e.g., MobileNet, EfficientNet-ES) replaces all depthwise Convs with sparse GroupNL models.
+As shown in Table IV, Sparse GroupNL-based CNNs achieve similar accuracy to standard lightweight CNN models on ImageNet-1K.
+As shown, GroupNL’s accuracy is 0.2% lower than MobileNet-V2’s and achieves the same accuracy as EfficientNet-ES.
+This effectively validates that sparse GroupNL can serve as a lightweight alternative to depthwise Conv.
+4.2.2 Evaluation on Corrupted Datasets
+TABLE V: Robustness on CIFAR-10-C.
+Method Noise Blur Weather Digital mAcc ($\uparrow$)
+ResNet-18 Vanilla CNN $52.60$ $72.78$ $84.55$ $80.70$ $72.58$
+MonoCNN $56.36$ $63.24$ $77.79$ $75.17$ $68.14$
+GhostNet $59.26$ $72.57$ $82.99$ $79.95$ $73.69$
+SineFM $53.77$ $68.60$ $84.63$ $80.94$ $71.99$
+GroupNL $57.55$ $72.96$ $85.25$ $80.68$ $74.11$
+ResNet-34 Vanilla CNN $53.72$ $71.70$ $85.08$ $82.27$ $73.19$
+MonoCNN $57.16$ $63.50$ $78.78$ $75.81$ $68.81$
+GhostNet $60.30$ $74.57$ $84.10$ $80.71$ $74.92$
+SineFM $59.20$ $71.72$ $86.11$ $81.83$ $74.72$
+GroupNL $59.84$ $73.64$ $86.01$ $81.26$ $75.19$
+ResNet-101 Vanilla CNN $53.97$ $73.05$ $86.02$ $81.90$ $74.78$
+MonoCNN $49.03$ $74.36$ $86.60$ $82.00$ $74.26$
+GhostNet $55.03$ $74.00$ $86.13$ $82.45$ $74.40$
+SineFM $58.33$ $72.08$ $85.08$ $81.44$ $74.23$
+GroupNL $57.90$ $71.22$ $86.45$ $81.17$ $75.04$
+VGG11 Vanilla CNN $65.37$ $75.45$ $82.06$ $80.48$ $75.84$
+MonoCNN $42.74$ $68.73$ $78.08$ $76.58$ $66.53$
+GhostNet $52.91$ $70.28$ $80.05$ $78.65$ $70.47$
+SineFM $62.31$ $72.46$ $82.46$ $82.61$ $74.96$
+GroupNL $64.10$ $77.24$ $83.46$ $83.87$ $77.17$
+TABLE VI: Robustness Performance Comparison on ImageNet-C.
+Model mCE ($\downarrow$) Corruption Error ($\downarrow$)
+Gauss. Noise Shot Noise Impulse Noise Defocus Blur Glass Blur Motion Blur Zoom Blur Snow Frost Fog Bright. Contrast Elastic Trans. Pixelate JPEG Comp. Speckle Noise Gauss. Blur Spatter Saturate
+Mobi-V2 64.109 73.044 73.420 73.202 72.645 78.768 66.340 71.765 70.056 65.514 58.210 37.570 61.305 61.230 76.440 52.273 64.376 68.692 50.543 42.685
+Mobi-V2-G 63.474 73.338 74.317 75.144 70.535 77.537 65.395 71.093 69.732 65.509 59.516 37.782 60.291 59.023 66.907 51.851 65.288 66.930 52.895 42.928
+Mobi-V3 58.501 67.779 68.460 66.505 66.762 74.732 58.622 65.558 62.274 60.467 54.441 33.461 58.942 56.282 70.918 44.757 58.332 62.887 44.492 37.376
+Mobi-V3-G 58.263 66.152 67.532 67.459 68.143 73.531 58.612 65.985 62.770 60.038 53.293 33.786 58.410 56.110 61.212 48.098 58.707 64.125 44.745 38.278
+Eff-Lite0 60.141 70.855 71.454 66.648 68.248 75.986 63.609 68.626 64.140 61.470 53.814 34.310 56.193 59.814 70.032 46.916 61.542 64.089 46.192 38.750
+Eff-Lite0-G 61.034 67.958 69.381 68.320 67.141 77.247 62.924 70.036 65.105 61.960 54.508 34.728 66.014 59.419 74.681 48.368 60.775 63.917 48.138 39.002
+Eff-ES 55.033 62.052 63.339 65.818 64.430 73.497 58.716 65.290 59.704 54.999 47.713 30.732 45.922 57.111 60.785 44.755 54.082 60.322 41.914 34.451
+Eff-ES-G 53.958 60.843 61.825 65.581 62.466 72.968 58.113 63.870 57.496 54.063 47.302 30.174 46.991 55.622 58.412 43.345 53.372 58.514 40.574 33.670
+Eff-V2-S 41.551 41.165 42.186 40.098 51.918 63.25 43.122 53.604 43.947 41.364 42.209 24.209 35.391 47.551 42.902 34.116 36.144 48.70 30.879 26.714
+Eff-V2-S-G 41.720 42.267 43.691 42.394 51.855 63.165 43.023 52.579 45.158 42.225 39.69 24.017 37.828 47.602 37.545 33.616 37.396 49.422 32.361 26.842
+Table V and Table VI show the performance comparison of different models for each method on the corrupted data.
+We observe that the GroupNL-based CNN models achieve the highest accuracy among alternatives to standard Conv.
+As shown in Table V, when ResNet-34 is used as the backbone, the accuracy of GroupNL is 0.47% higher than that of SineFM, 6.38% higher than that of MonoCNN, 0.27% higher than that of GhostNet, and even 2% higher than the accuracy of standard ResNet-34.
+When VGG11 is used as the backbone, the accuracy of GroupNL is 2.21% higher than that of SineFM, 10.64% higher than that of MonoCNN, 6.7% higher than that of GhostNet, and 1.33% higher than the accuracy of standard VGG11.
+There are two reasons for this: (i) The rules of the NLF can regularize the model and improve the generalization ability of the model.
+(ii) NLFs equipped with different hyperparameters can generate diverse feature maps and improve the knowledge expression ability of the model.
+Therefore, GroupNL achieves the highest accuracy.
+Sparse GroupNL also achieves considerable performance as a replacement for lightweight CNN models on ImageNet-C.
+As shown in Table VI, the models based on sparse GroupNL even achieve higher performance compared to the standard MobileNet-V2/V3, and EfficientNet-ES.
+This further validates that sparse GroupNL can serve as a lightweight alternative to depthwise Convs when dealing with corrupted datasets.
+TABLE VII: Robustness to Different Styles on Icons-50.
+Method Apple Facebook Google Samsung mAcc ($\uparrow$)
+ResNet-18 Vanilla CNN $93.66$ $87.64$ $83.20$ $80.28$ $86.19$
+MonoCNN $93.02$ $85.37$ $81.60$ $78.59$ $84.65$
+GhostNet $92.07$ $85.09$ $78.98$ $77.61$ $83.43$
+SineFM $94.84$ $88.72$ $84.96$ $82.27$ $87.70$
+GroupNL $94.52$ $91.05$ $86.27$ $84.36$ $89.05$
+ResNet-34 Vanilla CNN $92.77$ $86.26$ $83.81$ $82.44$ $86.32$
+MonoCNN $91.46$ $85.88$ $81.55$ $79.23$ $84.53$
+GhostNet $89.87$ $84.70$ $79.61$ $73.76$ $81.98$
+SineFM $94.74$ $88.72$ $84.88$ $81.46$ $87.45$
+GroupNL $95.19$ $90.98$ $85.93$ $84.36$ $89.12$
+ResNet-101 Vanilla CNN $93.64$ $87.70$ $83.64$ $83.13$ $87.03$
+MonoCNN $94.84$ $87.64$ $83.62$ $82.71$ $87.20$
+GhostNet $93.50$ $88.53$ $83.62$ $84.50$ $87.54$
+SineFM $90.25$ $83.65$ $79.69$ $74.67$ $82.06$
+GroupNL $93.98$ $90.15$ $85.51$ $83.89$ $88.38$
+VGG11 Vanilla CNN $92.54$ $86.68$ $83.67$ $80.34$ $85.81$
+MonoCNN $89.55$ $81.10$ $79.82$ $75.41$ $81.47$
+GhostNet $91.40$ $88.15$ $80.71$ $79.33$ $84.90$
+SineFM $94.74$ $89.87$ $85.59$ $82.17$ $88.09$
+GroupNL $94.23$ $91.05$ $85.30$ $82.13$ $88.18$
+4.2.3 Evaluation on Dataset under Different Styles
+We obtain conclusions similar to those for the corrupted data; the GroupNL-based CNN models achieve the highest accuracy under all backbones.
+As shown in Table VII, when ResNet-18 is used as the backbone, the accuracy of GroupNL is 1.35% higher than that of SineFM, 4.4% higher than that of MonoCNN, 5.62% higher than that of GhostNet, and even 2.86% higher than the accuracy of standard ResNet-18.
+when ResNet-101 is used as the backbone, the accuracy of GroupNL is 6.3% higher than that of SineFM, 1.18% higher than that of MonoCNN, 0.84% higher than that of GhostNet, and even 1.35% higher than the accuracy of standard ResNet-18.
+The reason for this is the same as that of our corrupted data; that is, (i) the nonlinear transformation function in GroupNL regularizes the model; (ii) GroupNL can generate diverse feature maps.
+Therefore, GroupNL has the highest accuracy.
+4.2.4 Training Acceleration Evaluation
+TABLE VIII: Comparison of CIFAR-10 Data Parallel Training Time in Different CNN methods in RTX 2080Ti GPUs.
+No. of GPUs LR BS Method Top-1 Acc (%) Training Time (h) Accel. Ratio
+[…]
+MonoCNN 94.97 9.2h +0%
+GhostNet 95.04 12.7h -37%
+SineFM 95.06 15.4h -66%
+GroupNL 95.51 8.3h +11%
+We verify the training acceleration of each CNN method based on ResNet-101. We observe that:
+(i) Compared to vanilla ResNet-101, GroupNL ResNet-101 achieves 11% speedup while achieving the higher accuracy, as shown in Table VIII.
+Moreover, since only two GPUs are used, the GroupNL mainly accelerates model training by reducing model calculation time.
+(ii) Although the GhostNet and SineFM have fewer parameters and computational complexity than vanilla CNN, the training time is longer, as shown in Table VIII.
+As we analyzed in Section 2.2, although cheap Conv and BN require less computation, they are slow to compute and optimize.
+Incorporating these operations into the model will significantly slow down the model’s training speed.
+We further compare the speed of GroupNL and vanilla CNN with different numbers of GPUs under DP and DDP.
+Training Evaluation in DP. For the multi-GPUs evaluation using DP, Table IX and Fig. 6 show the experimental settings and training results, respectively. We observe that:
+TABLE IX: Training Settings in DP/DDP Evaluation.
+Platform Parallelism No. of GPUs LR BS
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+With more GPUs, the GroupNL has a larger training speed-up ratio.
+As shown in Fig.
+6, when the number of GPUs is 4, GroupNL has the largest speed-up ratio compared to the vanilla CNN.
+This is because when the number of GPUs $N$ is large in DP, gradient communication $(N\!-\!1)\mathcal{G}$ for the main GPU and $\mathcal{G}$ for the rest (N1)(N\!-\!1) GPUs will increase when training a vanilla CNN.
+In contrast, the GroupNL with fewer gradients is less affected by the scale of GPUs.
+Therefore, when a large number of GPUs are used, GroupNL has a higher training speedup ratio.
+Figure 6: CIFAR-10 Training Acceleration Evaluation with Data Parallelism in RTX 2080Ti GPUs.
+Training Evaluation in DDP.
+For the multi-GPUs evaluation with DDP, Table IX and Fig.
+7 show the experimental settings and evaluated results of DDP training with AMP, respectively.
+We observe that: The DDP uses Ring-AllReduce to reduce gradient communication to $\frac{2(N-1)}{N}\mathcal{G}$ between GPUs.
+The GroupNL also reduces training time compared to vanilla CNNs, mainly due to fewer gradients.
+Figure 7: CIFAR-10 Training Acceleration Evaluation with DDP and AMP in RTX 2080Ti GPUs.
+Figure 8: ImageNet-1K Training Acceleration Evaluation with DDP (Parallelism ×8) in RTX 4090 GPUs.
+We also show the training speedup results for ImageNet-1K training of 8 GPUs.
+The GroupNL reduces training time by about 53% while achieving similar accuracy to the vanilla CNN.
+As shown in Fig 8.
+The $\times$8 DDP training vanilla CNN for 10 epochs is 3.32 hours with a Top-1 accuracy of 52.60%, and GroupNL is 1.87 hours with a Top-1 accuracy of 52.65%.
+For the $\times$8 DDP training vanilla CNN for 300 epochs is 151.93 hours with a Top-1 accuracy of 81.04%, and GroupNL is 72.11 hours with a Top-1 accuracy of 80.16%.
+Communication Cost & Epoch Time.
+We also evaluate the communication overhead and training time for 8-GPU DDP training of ResNet-101 and the corresponding GroupNL variant.
+With sizes of 44.55M and 26.37M gradients in models, respectively, the per-GPU communication cost and total 8-GPUs’ communication costs are 77.96M and 623.69M for ResNet-101, and 46.14M and 369.13M for GroupNL ResNet-101, respectively.
+Using the $\mathrm{epoch}=149$ training as an example, GroupNL ResNet-101 completes this epoch training in 1045.79s, with an average batch time of 0.827s, compared to 1754.07s and 1.398s for the vanilla version.
+This represents a 40% reduction in training time and a 41% reduction in gradients communication, highlighting the efficiency of GroupNL in acceleration.
+In general, GroupNL optimizes the model acceleration in terms of gradients communication and calculation time.
+4.3 Deployment and Resource Measurements
+4.3.1 Cloud-Device Transmission Cost
+We follow the cloud-device deployment pipeline in [42]: For PyTorch-CPU, we transmit well-trained PyTorch models; For ONNX Runtime, we convert PyTorch models to optimized ONNX models via gradient removal and BN fusion, and then transmit the ONNX models to the Raspberry Pi.
+TABLE X: Cloud-Device Transmission Model Size.
+Model PyTorch Size (Trained Model) ONNX Size
+MobileNet-V3 84.2MB 20.9MB
+MobileNet-V3-GroupNL 84.2MB 21.1MB
+EfficientNet-ES 62.6MB 20.7MB
+EfficientNet-ES-GroupNL 62.6MB 20.9MB
+ResNet-50 292.9MB 97.4MB
+ResNet-50-GroupNL 199.1MB 66.3MB
+ResNet-101 510.7MB 169.8MB
+ResNet-101-GroupNL 302.6MB 100.7MB
+As shown in Table X, GroupNL maintains nearly identical sizes for lightweight CNNs like MobileNet-V3 and EfficientNet-ES in both PyTorch and ONNX formats.
+Also, it achieves substantial size reductions in larger-size CNNs, decreasing ResNet-50 by approximately 32% and ResNet-101 by about 41%.
+This demonstrates GroupNL’s advantage for efficient cloud-device transmission.
+4.3.2 Resource Evaluations in On-Device Inference
+We infer 100 images randomly selected from ImageNet-1K validation datasets on a resource-constrained Raspberry Pi to calculate the inference time, average Frames Per Second (FPS), memory usage, average power, and energy consumption.
+The average FPS computation follows the setting of this work [44].
+TABLE XI: On-Device Evaluation with PyTorch-CPU.
+Model Time (Sec) Avg. FPS ($\uparrow$) Avg. Resource Usage Acc (%)
+Mem. (MB) Power (W) Energy (Wh)
+Mobi-V3 52.9 1.89 180.0 6.144 0.0973 86
+Mobi-V3-G 43.5 2.30 181.0 6.034 0.0805 84
+Eff-ES 122.0 0.82 187.9 6.101 0.2102 88
+Eff-ES-G 95.1 1.05 206.1 5.866 0.1629 88
+RS-50 106.7 0.94 269.4 5.171 0.1623 91
+RS-50-G 86.7 1.15 228.8 5.452 0.1363 92
+RS-101 192.7 0.52 340.0 5.125 0.2819 94
+RS-101-G 146.4 0.68 262.0 5.518 0.2345 94
+PyTorch-CPU Backend.
+As shown in Table XI, GroupNL significantly boosts inference efficiency on the Raspberry Pi.
+For example, in EfficientNet-ES, it reduces inference time and energy consumption by up to 22% and 22.5%, respectively.
+It also maintains competitive accuracy and reduces memory overhead in larger models like ResNet-50 and ResNet-101.
+As shown in Fig.
+9, for power consumption (left), GroupNL yields shorter inference times and lower peak power than its counterparts; For energy consumption (right), the shorter inference time of GroupNL results in lower total energy.
+Figure 9: Power and Energy Consumption in PyTorch-CPU.
+TABLE XII: On-Device Evaluation with ONNX Runtime.
+Model Time (Sec) Avg. FPS ($\uparrow$) Avg. Resource Usage Acc (%)
+Mem. (MB) Power (W) Energy (Wh)
+Mobi-V3 13.0 7.67 174.0 6.347 0.0277 86
+Mobi-V3-G 15.8 6.35 178.0 6.290 0.0330 84
+Eff-ES 29.1 3.44 176.0 6.698 0.0588 88
+Eff-ES-G 32.8 3.04 183.0 6.671 0.0660 88
+RS-50 44.5 2.25 270.9 6.772 0.0922 91
+RS-50-G 35.4 2.83 236.0 6.795 0.0719 92
+RS-101 72.3 1.38 343.9 6.841 0.1497 94
+RS-101-G 55.9 1.79 272.9 6.766 0.1152 94
+ONNX Runtime Backend.
+As shown in Table XII, GroupNL improves inference efficiency for larger models while showing limited gains on lightweight CNNs when using ONNX Runtime.
+For instance, in ResNet-101, it reduces inference time and energy consumption by approximately 22.7% and 23%, respectively.
+However, in lightweight CNNs such as EfficientNet-ES, it currently incurs a slight increase in latency and energy use.
+This is because the converted GroupNL ONNX model does not yet support efficient tensor manipulation operators, e.g., one $\tt{torch.repeat}$ is converted into multiple $\tt{onnx.tile}$ and $\tt{onnx.expand}$ operations, introducing overhead that diminishes its advantage compared with depthwise Conv.
+A similar power and energy consumption is reported in Fig.
+10.
+Figure 10: Power and Energy Consumption in ONNX Runtime.
+4.4 Ablation Studies
+4.4.1 Selection of Nonlinear Functions
 Choosing appropriate nonlinear transformations and selecting appropriate hyperparameters for the NLF is highly related to the performance of the GroupNL model.
@@
 TABLE 13: Nonlinear transformation functions.
-Nonlinear Transformation Formulation Hyperparameters
+Nonlinear Formulation Hyperparameters
 Sinusoidal $\mathrm{sin}(\omega(x+\phi))$ $\omega,\phi\in\mathbb{R}$
@@
 Laplace $\frac{\epsilon}{2}e^{-(\epsilon x)}$ $\epsilon\in\mathbb{R}$
-In this case, we can observe that the GroupNL model with Sinusoidal function outperforms the other three NLFs and also has a similar performance to the vanilla CNN model.
-As a result, we can generate feature maps by the Sinusoidal function as the NLF in subsequent experiments.
-Figure 4: Comparison of different NLFs in GroupNL ResNet-18 on the CIFAR-10 dataset.
+In this case, we can observe that the GroupNL model with the Sinusoidal function outperforms the other three NLFs and also has a similar performance to the vanilla CNN model.
+As a result, in our experiments, we use the Sinusoidal function as the NLF to generate feature maps.
+We offer a conjecture: due to the periodicity of the Sinusoidal function and its different frequencies and phase hyperparameters, it can generate a richer variety of “similar but phase-different patterns” compared to other functions, which meets the requirement of GroupNL to generate similar but diverse feature maps.
+The reasons why other functions are worse may be: Monomial only performs power transformations and does not generate new spatial patterns, resulting in insufficient diversity; Gaussian function has too strong locality, with values only near the function center, and other values are suppressed to 0, which cannot guarantee the similarity of the generated images; Laplace function is similar to Gaussian.
+Figure 11: Comparison of different NLFs in GroupNL and vanilla ResNet-18 on the CIFAR-10 dataset.
+4.4.2 Selection of Hyperparameters and Coefficients
 Analysis of Sparse Seed Filters in GroupNL: As described in Section 3.2.1, we designate that the seed filters ${\bm{W}}_{\mathrm{seed}}$ in basic structure (e.g., ResNet-18) utilize the standard convolution, and the seed filters in bottleneck structure (e.g., ResNet-101) are sparse with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$.
 The sparse seed filter experimental results are shown in Table 14.
-Applying sparsity to the seeded convolution of ResNet-18 can introduce an extremely low number of parameters of 0.05M and just approximately 7M FLOPs.
+Applying sparsity to the seeded convolution of ResNet-18 can introduce an extremely low number of parameters of 0.05M and approximately 7M FLOPs.
 However, performance is severely degraded.
-As a comparison, the ResNet-101 with sparsity can maintain the performances with the lowest number of parameters and FLOPs.
+As a comparison, the ResNet-101 with sparsity can maintain the performance with the lowest number of parameters and FLOPs.
 Therefore, the sparse seed filters are not used in the basic structure.
-TABLE 3: The ablation studies of w/ and w/o sparse seed filters and different reduction ratio $r$ in GroupNL.
-Model $g$ $r$Params (M) #FLOPs (M) Top-1 Acc (%)
+Additionally, Sparse GroupNL modules are designated to replace the depthwise convolutions in lightweight CNNs [34, 17, 8, 36, 7, 37] to maintain the compactness of these models in terms of number of parameters and FLOPs.
+TABLE 14: Ablation Studies of w/ and w/o Sparse Seed Filters and Reduction Ratio $r$ in GroupNL.
+Model $g$ $r$ #Params (M) #FLOPs (M) Top-1 Acc (%)
 ResNet-18 (w/o sparse) 4 2 5.6 279 95.12
@@
 Hyperparameters and Coefficients of GroupNL: We designate that the hyperparameters of NLFs are randomly initialized and non-learnable during model training.
-This trick reduces the number of gradients for optimization in training while preventing the NLFs lose the data-agnostic properties.
+This trick reduces the number of gradients for optimization in training while preventing the NLFs from losing the data-agnostic properties.
 The Sinusoidal function has two hyperparameters $\omega$ and $\phi$, we randomly choose them via $\tt{torch.rand}$ between $[1,2]$ and $[1,5]$, respectively.
@@
 Besides, for the ResNet-101, the number of parameters and FLOPs are very similar for different $r$ settings.
-This is due to the bottleneck structure in ResNet-101 using the sparse seed filters with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$.
+This is due to the bottleneck structure in ResNet-101 using the sparse seed filters with convolutional groups of $\xi=\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$.
 For example, assuming a $k^{2}$ convolution with same input and output channels $c_{\mathrm{in}}\!=\!c_{\mathrm{out}}\!=\!512$ and feature maps size $wh$ in the bottleneck.
-In the case of $r=2$, the number of parameters is $\mathcal{P}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}}{256}=512k^{2}$, and the FLOPs is $\mathcal{F}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}wh}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}wh}{256}=512k^{2}wh$.
+In the case of r=2r\!=\!2, the number of parameters is $\mathcal{P}_{r2}\!=\!c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}\frac{1}{\xi}\!=\!512\cdot 256k^{2}\frac{1}{256}\!=\!512k^{2}$, and the FLOPs is $\mathcal{F}_{r2}\!=\!whc_{\mathrm{in}}c_{\mathrm{seed}}k^{2}\frac{1}{\xi}\!=\!wh\cdot 512\cdot 256k^{2}\frac{1}{256}\!=\!512k^{2}wh$.
 In the cases of $r\!=\!\{4,16,32,\cdots,256,512\}$, the results are also $\mathcal{P}_{r}\!=\!512k^{2}$ and $\mathcal{F}_{r}\!=\!512k^{2}wh$ because of the fraction $c_{\mathrm{seed}}$ are reduced in the case of using the greatest common divisor grouping seed filters.
 For the $g$, on the other hand, because the parameters and computations depend on the seed filters, different $g$ will not affect the FLOPs and the number of parameters.
-TABLE 4: The ablation studies of different splitting groups $g$ in GroupNL CNN models.
-Model $r$ $g$Params (M) #FLOPs (M) Top-1 Acc (%)
+TABLE 15: Ablation Studies of Groups $g$ in GroupNL.
+$r$ $g$ #Params (M) #FLOPs (M) Top-1 Acc (%)
 ResNet-18 2 4 5.6 279 95.12
@@
 16 18.5 1159 95.25
-3.4 Resource and Training Acceleration Analysis
-For cloud-assisted CNN models, the multi-GPUs training on cloud servers is mainly based on data parallelism.
-In practical multi-GPU training, the bandwidth of inter-GPU memory data transfer is an important constraint on the training speed compared to the powerful and fast computation kernels.
-In the classic data parallelism, e.g., $\tt{torch.nn.DataParallel}$, the data exchange between different GPUs in forward/backward of the model generally in three stages: (i) Gather outputs (or logits) and compute loss gradients on host GPU; (ii) Scatter gradients to multi-GPUs; (iii) Reduce gradients to host GPUs.
-For the distributed data parallelism, e.g., $\tt{torch.nn.parallel.DistributedDataParallel}$, only one reduction of gradients between GPUs in each backward.
-So, the number of gradients can be further optimized in terms of inter-GPU scatters and reductions in different data parallelism.
-Assuming each trainable parameter in a CNN model maintains a gradient, the number of gradients $\mathcal{G}$ corresponds to the number of trainable parameters $\mathcal{P}$.
-At the same time, designing libraries-agnostic and task-specific CNN methods is the minimal change in cloud-assisted architecture for generalization.
-Based on this premise, we analyze the theoretical gradient exchange in multi-GPUs memory of vanilla CNN, GhostNet, SineFM, and GroupNL in training.
-Ignoring the bias, for a vanilla convolutional layer, the number of parameters and gradients can be described as follows:
-$\begin{array}[]{l}{\mathcal{P}_{\mathrm{vanilla\ cnn}}=\mathcal{G}}_{\mathrm{vanilla\ cnn}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\ .\end{array}$ (13)
-For a GhostNet layer, the number of parameters and gradients from the seed filters and $d\times d$ DWConv are:
-${\mathcal{P}_{\mathrm{ghostnet}}=\mathcal{G}}_{\mathrm{ghostnet}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+c_{\mathrm{gen}}d^{2}$ (14)
-$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!-\!\frac{1}{r})d^{2}\ ,$
-and the number of parameters and gradients for the SineFM layer from the seed filters, $1\times 1$ DWConv, and batch normalization are:
-${\mathcal{P}_{\mathrm{sinefm}}=\mathcal{G}}_{\mathrm{sinefm}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+tc_{\mathrm{gen}}\cdot 1^{2}+2tc_{\mathrm{seed}}$ (15)
-$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+tc_{\mathrm{out}}(1\!-\!\frac{1}{r})+2tc_{\mathrm{out}}/r$
-$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!+\!\frac{1}{r})t\ .$
-The proposed GroupNL layer only contains the learnable standard seed filters and learnable sparse seed filters, and the number of parameters and gradients are shown in Eq. 16 and Eq. 17, respectively:
+4.4.3 Module-Level Profiling
+We conduct a module-level profiling of GroupNL Conv on Raspberry Pi, based on PyTorch-CPU backend.
+For Conv modules, the number of input and output channels is 512, the input feature maps size is (512, 32, 32), kernel size is (3, 3), stride is 2, and padding is 1, so the output feature maps size is (512, 16, 16).
+The FPS reports the average Frames Per Second in 100-times inference with randomly initialized inputs.
 […]
-[…]
-Noticeably, the relation is $\mathcal{G}_{\mathrm{groupnl}}^{\mathrm{sparse}}\leq\mathcal{G}_{\mathrm{groupnl}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r<\mathrm{min}\{{\mathcal{G}}_{\mathrm{cnn}},{\mathcal{G}}_{\mathrm{ghostnet}},{\mathcal{G}}_{\mathrm{sinefm}}\}$, i.e., the number of gradients for optimization in GroupNL is minimal compared to the rest CNN design methods of ${\mathcal{G}}_{\mathrm{cnn}}$, ${\mathcal{G}}_{\mathrm{ghostnet}}$, and ${\mathcal{G}}_{\mathrm{sinefm}}$.
-The sparse GroupNL also significantly reduces the gradient exchange by a factor of $\frac{1}{r\cdot\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{out}}/r)}$ compared to the ${\mathcal{G}}_{\mathrm{cnn}}$.
-As the number of GPUs scales, this scatter/reduction gap compared to vanilla CNNs also grows larger thus accelerating the training.
-It is notable that, it seems that GhostNet and SineFM also have a small number of gradients with ${{\mathcal{G}}_{\mathrm{ghostnet}}<\mathcal{G}}_{\mathrm{cnn}}$ and ${\mathcal{G}}_{\mathrm{sinefm}}<{\mathcal{G}}_{\mathrm{cnn}}$, can also accelerate training in theory.
-However, the evaluation in Table I shows that the GhostNet and SineFM are even slower than the vanilla CNN.
-For this phenomenon, we consider this to be due to the introduction of more computational nodes involved in the backward updating, such as DWConv and Norm, which leads to slower training instead.
-Our proposed GroupNL, on the other hand, does not introduce new computational nodes through in-place operations such as $\tt{torch.cat}$ and $\tt{torch.split}$, and copy operations such as $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, thus ensures acceleration of training.
-3.5 Discussion
-As the aforementioned analysis, compared with the GhostNet [18] and SineFM [19], the proposed GroupNL-based CNN design method has two highlights:
-(i) More diverse feature maps.
-Given the seed filters and the number of output channels of the layer, by grouping the seed filters, more different hyperparameter settings of data-agnostic NLFs can be used to generate diverse features.
-(ii) More training efficient.
-GroupNL generates the number of feature maps without extra computational nodes like DWConv and Norm, effectively reducing the number of gradients exchanging in the multi-GPUs training.
-4 Experiments
-This section introduces our experimental setup, including the datasets, baselines, and implementation details.
-Then, we empirically compare accuracy, network traffic, FLOPs, and training speed on multiple benchmarks.
-4.1 Experimental Setup
-4.1.1 Datasets
-We evaluate the effectiveness of the proposed method using five widely used datasets.
-CIFAR-10, GTSRB, and ImageNet-1K are standard datasets.
-CIFAR-10-C and Icons-50 are used as corrupted datasets.
-Among them, noise is introduced in the CIFAR-10-C dataset, and the training data and test data styles of Icons-50 are quite different.
-CIFAR-10 [35] is a multi-class natural object dataset for image classification.
-It consists of 50,000 training images and 10,000 test images in 10 categories, each with a resolution of $32\times 32$ pixels.
-GTSRB [36] is a German traffic sign dataset for traffic sign recognition.
-It comprises 39,209 training images and 12,630 test images from 43 categories.
-The images in GTSRB have varying light conditions and rich backgrounds.
-We set the resolution of all images to $32\times 32$ pixels.
-CIFAR-10-C [37] is a test dataset after using synthetic common perturbations and noise corruptions on the CIFAR-10 test set.
-It consists of 10,000 test images of 19 types of damage in 4 categories.
-The resolution of each image is $32\times 32$ pixels.
-Icons-50 [37] consists of 10,000 images in 50 categories collected from different companies.
-We set the resolution of all images to $32\times 32$ pixels.
-For training, the data of one company is retained as test data, and the data of the rest of the companies is used as training data.
-Please refer to [5] for more explanation of the CIFAR-10-C and Icons-50 datasets.
-ImageNet [38] is a large-scale classification dataset to verify computer vision performance.
-ImageNet-1K is a subset of the ILSVRC ImageNet, which contains 1.28 billion training images and 50K validation images from 1K different classes.
-The resolution of each image is $224\times 224$ pixels.
-4.1.2 Baseline CNNs
-We compare the following nonlinear transformation and feature maps generated-based methods in CNNs.
-MonoCNN [5]: In MonoCNN, the convolutional layer is specified with only a few seed filter parameters updated during model training.
-The remaining filters in this layer are generated based on the seed filters and monomial function.
-The hyperparameters of the monomial function are randomly generated and non-learnable in training.
-GhostNet [18]: In GhostNet, a few filters are first used to generate some feature maps.
-Then, utilizing $d\times d$ DWConv operations to generate the remaining feature maps based on these generated feature maps.
-SineFM [19]: In the SineFM-based CNN model, a small number of seed convolutional filters are first specified in the convolutional layer.
-Then, seed feature maps are generated based on the seed filters and input data.
-Subsequently, a set of expanding feature maps is generated based on the seed feature map, $t$-sinusoidal functions, and a $t$-sets of BatchNorm.
-Finally, utilizing the $1\times 1$ DWConv to align the number of expanding features to the generated features.
-To ensure a comprehensive comparison, all baseline methods are based on the following CNN models: ResNet-18/-34/-101 [2] and VGG11 [1].
-To make the effect of training acceleration more clear, we evaluate the training speed of the larger model, i.e., the ResNet-101.
-4.1.3 Implementation Details
-The implementation details include training platforms & libraries and experimental settings.
-Training Platforms and Libraries: The experiments are performed on the servers with NVIDIA RTX 2080Ti 11GB and RTX 4090 GPUs 24GB.
-We implement the models in PyTorch 1.12 for the NVIDIA platforms with CUDA 11.1.
-Hyperparameters Settings of Baseline CNNs: For the hyperparameters setting, we follow the suggestion from the original papers [5, 18, 19].
-The specific description of model hyperparameters is shown in Table 5.
-(i) MonoCNN [5] uses the monomial function as a filter generating function and samples a random and continuous exponent $\eta$ between $[1,7]$, and the exp_factor is $\{2,4,8,16\}$;
-(ii) GhostNet [18] uses $3\times 3$ grouped convolution with $d=3$ as the cheap operation for generating the feature maps and sets the reduction ratio to 2;
-(iii) SineFM [19] uses a sinusoidal function as the feature map generator function with random hyperparameters, the real numbers of period $\omega$ and shift $\phi$, which ranges are between $[1,2]$ and $[1,5]$, respectively.
-The number of the NLFs with different hyperparameter settings in SineFM is $t=5$, which means five times expanding the seed features as described in Eq.
-7.
-Experimental Settings: For the settings of training in CIFAR-10/GTSRB/Icons-50, the learning rate is 0.1, and the batch size is 128, with the SGD optimizer and cosine annealing scheduler.
-For the training in ImageNet-1K, the learning rate is 0.001, and the batch size is 1024, with the Adam optimizer and cosine annealing scheduler.
-In addition, the training evaluation is compared for the larger model, i.e., ResNet-101 in different parallel technologies in PyTorch library API [39], including the Data Parallelism (DP) with $\tt{torch.nn.DataParallel}$, the Distributed Data Parallelism (DDP) [40] with $\tt{torch.nn.parallel.DistributedDataParallel}$, and the Automatic Mixed Precision (AMP) [41] with $\tt{torch.cuda.amp}$.
-We also follow the linear scaling rule [42] to scale the number of parallelism in training, which is scaling the learning rate by the batch size.
-In addition, we also set up the same scale of seed feature map numbers in GhostNet, SineFM, and GroupNL with $r=2$.
-TABLE 5: Hyperparameter setting of CNNs methods.
-Method Hyperparameters Settings
-Reduction Coefficient & Random Range
-MonoCNN $\mathrm{exp\_factor}=\{2,4,8,16\}$ $\eta\in[1,7]$
-GhostNet $r=2$ $d=3$
-SineFM $r=2$ $t=5,\quad\omega\in[1,2],\quad\phi\in[1,5]$
-GroupNL $r=2$ $g=4,\quad\omega\in[1,2],\quad\phi\in[1,5]$
-4.2 Experimental Results
-4.2.1 Reduction of Network Traffic and FLOPs
-Figure 5: Number of learnable parameters for different models with different methods.
-Figure 6: FLOPs for different models with different methods.
-Since the non-learnable hyperparameters of NLFs are randomly initialized and remain unchanged, these parameters can be saved and reproduced with a random number.
-Therefore, the number of learnable parameters in the CNN model can be approximated to the number of model parameters that need to be transmitted when deploying the model on cloud-assisted IoT devices.
-Fig.
-5 and Fig.
-6 show the number of learnable parameters and FLOPs of different CNN models for different methods.
-Without using sparse technology, the number of learnable parameters of GhostNet, SineFM, and GroupNL is similar, which is much less than the number of learnable parameters of vanilla CNN.
-As shown in Fig.
-5, When based on VGG11, ResNet-18, and ResNet-34 architectures, GhostNet, SineFM, and GroupNL have a similar number of learnable parameters.
-When based on ResNet-101 architecture, GroupNL has the least amount of learnable parameters due to the user of sparse technology.
-As shown in Fig.
-6, in terms of FLOPs, there are similar findings as in the amount of learnable parameters.
-4.2.2 Robustness
-This section evaluates the robustness of the GroupNL-based CNN on three types of datasets, including standard datasets (i.e., CIFAR-10 and GTSRB), the corrupted dataset (i.e., CIFAR-10-C), and data under different styles (i.e., Icons50).
-1) Standard Data.
-TABLE 6: Comparison of image classification results on CIFAR-10 and GTSRB datasets.
-Methods CIFAR-10 Top-1 Acc (%) GTSRB Top-1 Acc (%)
-ResNet-18 ResNet-34 ResNet-101 VGG11 ResNet-18
-Vanilla CNN 95.04 95.24 95.48 91.89 97.89
-MonoCNN 92.05 92.26 94.97 88.84 97.66
-GhostNet 93.36 93.90 95.04 92.65 95.76
-SineFM 94.36 94.77 95.06 91.27 98.18
-GroupNL 95.12 95.45 95.51 91.62 98.31
-Table 6 shows the comparison of image classification results of different CNN models on CIFAR-10 and GTSRB datasets under different methods. We observe that:
-(i) The accuracy of the GroupNL-based CNN exceeds that of all alternatives to the standard CNN model.
-As shown, when using ResNet-34 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.68% higher than SineFM’s, 1.55% higher than GhostNet’s, and 3.19% higher than MonoCNN’s.
-When using VGG11 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.35% higher than SineFM’s.
-The main reason is that GroupNL generates diversified feature maps using the NLFs with different hyperparameter configurations.
-In contrast, SineFM generates feature maps using a single NLF and is unable to generate diverse feature maps.
-GhostNet first generates some feature maps using standard convolution operations and then uses cheap linear operations to generate other feature maps based on these generated feature maps, which limits the diversity of the feature maps.
-Benefiting from its ability to generate diverse feature maps, GroupNL achieves the highest accuracy among many alternatives.
-(ii) The GroupNL-based CNN models outperform the vanilla CNN models in most cases.
-As shown in Table 6, GroupNL performs slightly worse than the vanilla CNN only when using VGG11 as the backbone.
-When using ResNet-18/-34/-101, the accuracy of GroupNL exceeds that of the vanilla CNN.
-This is mainly due to the ability of GroupNL to generate diverse feature maps.
-2) Corrupted Data.
-TABLE 7: Robustness to common observable corruptions. Our results are highlighted with shading.
-Method Noise Blur Weather Digital mean
-Vanilla CNN $52.60$ $72.78$ $84.55$ $80.70$ $72.58$
-MonoCNN $56.36$ $63.24$ $77.79$ $75.17$ $68.14$
-GhostNet $59.26$ $72.57$ $82.99$ $79.95$ $73.69$
-SineFM $53.77$ $68.60$ $84.63$ $80.94$ $71.99$
-GroupNL $57.55$ $72.96$ $85.25$ $80.68$ $74.11$
-Method Noise Blur Weather Digital mean
-Vanilla CNN $53.72$ $71.70$ $85.08$ $82.27$ $73.19$
-MonoCNN $57.16$ $63.50$ $78.78$ $75.81$ $68.81$
-GhostNet $60.30$ $74.57$ $84.10$ $80.71$ $74.92$
-SineFM $59.20$ $71.72$ $86.11$ $81.83$ $74.72$
-GroupNL $59.84$ $73.64$ $86.01$ $81.26$ $75.19$
-Method Noise Blur Weather Digital mean
-Vanilla CNN $53.97$ $73.05$ $86.02$ $81.90$ $74.78$
-MonoCNN $49.03$ $74.36$ $86.60$ $82.00$ $74.26$
-GhostNet $55.03$ $74.00$ $86.13$ $82.45$ $74.40$
-SineFM $58.33$ $72.08$ $85.08$ $81.44$ $74.23$
-GroupNL $57.90$ $71.22$ $86.45$ $81.17$ $75.04$
-Method Noise Blur Weather Digital mean
-Vanilla CNN $65.37$ $75.45$ $82.06$ $80.48$ $75.84$
-MonoCNN $42.74$ $68.73$ $78.08$ $76.58$ $66.53$
-GhostNet $52.91$ $70.28$ $80.05$ $78.65$ $70.47$
-SineFM $62.31$ $72.46$ $82.46$ $82.61$ $74.96$
-GroupNL $64.10$ $77.24$ $83.46$ $83.87$ $77.17$
-Table 7 shows the performance comparison of different models for each method on the corrupted data. We observe that:
-The GroupNL-based CNN models achieve the highest accuracy.
-As shown in Table 7, when ResNet-34 is used as the backbone, the accuracy of GroupNL is 0.47% higher than that of SineFM, 6.38% higher than that of MonoCNN, 0.27% higher than that of GhostNet, and even 2% higher than the accuracy of standard ResNet-34.
-When VGG11 is used as the backbone, the accuracy of GroupNL is 2.21% higher than that of SineFM, 10.64% higher than that of MonoCNN, 6.7% higher than that of GhostNet, and 1.33% higher than the accuracy of standard VGG11.
-There are three reasons for this: (i) GroupNL sets some model parameters to be non-learnable, so that the model parameters are not completely obtained through training by inputting training data, thus making the model have stronger generalization ability.
-(ii) The rules of the NLF can regularize the model and improve the generalization ability of the model.
-(iii) NLFs equipped with different hyperparameters can generate diverse feature maps and improve the knowledge expression ability of the model.
-Therefore, GroupNL achieves the highest accuracy.
-The model’s ability to generate diverse feature maps has a great impact on the robustness of the model.
-As shown, with ResNet-18 as the backbone, GroupNL’s accuracy is 2.12% higher than SineFM’s.
-With ResNet-34 as the backbone, GroupNL’s accuracy is 0.47% higher than SineFM’s.
-With ResNet-101 as the backbone, GroupNL’s accuracy is 0.81% higher than SineFM’s.
-With the same number of parameters and computational complexity, the difference between GroupNL and SineFM is that the former can generate a variety of feature maps.
-This shows that increasing the diversity of feature maps can effectively improve the robustness of the CNN model.
-TABLE 8: Robustness to different styles. Our results are highlighted with shading.
-Method Apple Facebook Google Samsung mean
-Vanilla CNN $93.66$ $87.64$ $83.20$ $80.28$ $86.19$
-MonoCNN $93.02$ $85.37$ $81.60$ $78.59$ $84.65$
-GhostNet $92.07$ $85.09$ $78.98$ $77.61$ $83.43$
-SineFM $94.84$ $88.72$ $84.96$ $82.27$ $87.70$
-GroupNL $94.52$ $91.05$ $86.27$ $84.36$ $89.05$
-Method Apple Facebook Google Samsung mean
-Vanilla CNN $92.77$ $86.26$ $83.81$ $82.44$ $86.32$
-MonoCNN $91.46$ $85.88$ $81.55$ $79.23$ $84.53$
-GhostNet $89.87$ $84.70$ $79.61$ $73.76$ $81.98$
-SineFM $94.74$ $88.72$ $84.88$ $81.46$ $87.45$
-GroupNL $95.19$ $90.98$ $85.93$ $84.36$ $89.12$
-Method Apple Facebook Google Samsung mean
-Vanilla CNN $93.64$ $87.70$ $83.64$ $83.13$ $87.03$
-MonoCNN $94.84$ $87.64$ $83.62$ $82.71$ $87.20$
-GhostNet $93.50$ $88.53$ $83.62$ $84.50$ $87.54$
-SineFM $90.25$ $83.65$ $79.69$ $74.67$ $82.06$
-GroupNL $93.98$ $90.15$ $85.51$ $83.89$ $88.38$
-Method Apple Facebook Google Samsung mean
-Vanilla CNN $92.54$ $86.68$ $83.67$ $80.34$ $85.81$
-MonoCNN $89.55$ $81.10$ $79.82$ $75.41$ $81.47$
-GhostNet $91.40$ $88.15$ $80.71$ $79.33$ $84.90$
-SineFM $94.74$ $89.87$ $85.59$ $82.17$ $88.09$
-GroupNL $94.23$ $91.05$ $85.30$ $82.13$ $88.18$
-3) Data under different styles: We obtain conclusions similar to those for the corrupted data; the GroupNL-based CNN models achieve the highest accuracy under all backbones.
-As shown in Table 8, when ResNet-18 is used as the backbone, the accuracy of GroupNL is 1.35% higher than that of SineFM, 4.4% higher than that of MonoCNN, 5.62% higher than that of GhostNet, and even 2.86% higher than the accuracy of standard ResNet-18.
-when ResNet-101 is used as the backbone, the accuracy of GroupNL is 6.3% higher than that of SineFM, 1.18% higher than that of MonoCNN, 0.84% higher than that of GhostNet, and even 1.35% higher than the accuracy of standard ResNet-18.
-The reason for this is the same as that of our corrupted data; that is, (i) GroupNL makes the model parameters not completely updated by the input training data; (ii) the nonlinear transformation function in GroupNL regularizes the model; (iii) GroupNL can generate diverse feature maps.
-Therefore, GroupNL has the highest accuracy.
-4.2.3 Training Acceleration
-TABLE 9: Comparison of CIFAR-10 Training Time in Different CNN methods in RTX 2080Ti GPUs.
-Tech. Parallelism No. of GPUs LR BS Method Top-1 Acc (%) Training Time Accelerated Ratio
-Data Parallel $\times$1 2-GPUs 0.1 64$\cdot$2 Vanilla CNN 95.48 9.3 Hours -
-MonoCNN 94.97 9.2 Hours +0%
-GhostNet 95.04 12.7 Hours -37%
-SineFM 95.06 15.4 Hours -66%
-GroupNL 95.51 8.3 Hours +11%
-We verify the training acceleration of each CNN method based on ResNet-101. We observe that:
-(i) Compared to vanilla ResNet-101, GroupNL-based ResNet-101 achieves 11% speedup while achieving the higher accuracy, as shown in Table 9.
-This is because the GroupNL method uses the NLFs to generate feature maps, reducing the number of operations based on sliding windows to generate feature maps.
-The FLOPs of generating feature maps based on nonlinear transformation are significantly less than those of generating feature maps based on sliding windows (as shown in Fig.
-6).
-Moreover, since only two GPUs are used, the GroupNL mainly accelerates model training by reducing model calculation time.
-(ii) Although the GhostNet and SineFM have fewer parameters and computational complexity than vanilla CNN, the training time is longer, as shown in Table 9.
-As we analyzed in Section 2.1, although cheap linear transformations and normalizations require less computation, they are slow to compute and optimize.
-Incorporating these operations into the model will significantly slow down the model’s training speed.
-We then further compare the speedup of GroupNL and vanilla CNN with different numbers of GPUs under Data Parallel (DP) and Distributed Data Parallel (DDP).
-1) Training Evaluation in DP. For the multi-GPUs evaluation using DP technology, Table 10 and Fig. 7 show the experimental settings and training results, respectively. We observe that:
-TABLE 10: Training Settings in DP Evaluation.
-Platform Parallelism No. of GPUs LR No. of BS
-RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
-$\times$1 2-GPUs 0.1 64$\cdot$2
-$\times$2 2-GPUs 0.2 128$\cdot$2
-$\times$3 3-GPUs 0.3 128$\cdot$3
-$\times$4 4-GPUs 0.4 128$\cdot$4
-With more GPUs, the GroupNL has a larger training speed-up ratio.
-As shown in Fig.
-7, when the number of GPUs is 4, GroupNL has the largest speed-up ratio compared to the vanilla CNN.
-This is because when the number of GPUs is large in DP, the parameter transmission time between multiple GPUs will increase significantly when training vanilla CNN.
-In contrast, the non-trainable parameters in the GroupNL do not need to be transmitted and are less affected by the parameter transmission between multiple GPUs.
-Therefore, when a large number of GPUs are used, GroupNL has a higher training speedup ratio.
-Figure 7: CIFAR-10 Training Acceleration Evaluation with Data Parallelism in RTX 2080Ti GPUs.
-2) Training Evaluation in DDP.
-For the multi-GPUs evaluation with DDP technology, Table 11 and Fig.
-8 show the experimental settings and evaluated results of DDP training with AMP, respectively.
-We observe that: Since DDP uses Ring-AllReduce to reduce parameter transmission between multiple GPUs, the GroupNL reduces training time compared to the vanilla CNN mainly because it reduces the computational workload of the CNN model.
-This shows that the GroupNL can effectively speed up the training of the model by using nonlinear transformation to generate feature maps to reduce the use of sliding windows to generate feature maps, and also makes the GroupNL-based model more suitable for deployment on IoT devices.
-TABLE 11: Training Settings in DDP Evaluation.
-Platform Parallelism No. of GPUs LR No. of BS
-RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
-$\times$2 2-GPUs 0.2 128$\cdot$2
-$\times$4 2-GPUs 0.4 256$\cdot$2
-$\times$8 4-GPUs 0.8 256$\cdot$4
-RTX 4090 $\times$8 8-GPUs 0.001 128$\cdot$8
-Figure 8: CIFAR-10 Training Acceleration Evaluation with DDP and AMP in RTX 2080Ti GPUs.
-Figure 9: ImageNet-1K Training Acceleration Evaluation with DDP (Parallelism ×8) in RTX 4090 GPUs.
-We also show the training speedup results for ImageNet-1K training of 8 GPUs.
-The GroupNL reduces training time by about 53% while achieving similar accuracy to the vanilla CNN.
-As shown in Fig 9.
-The $\times$8 DDP training vanilla CNN for 10 epochs is 3.32 hours with a Top-1 accuracy of 52.60%, and GroupNL is 1.87 hours with a Top-1 accuracy of 52.65%.
-For the $\times$8 DDP training vanilla CNN for 300 epochs is 151.93 hours with a Top-1 accuracy of 81.04%, and GroupNL is 72.11 hours with a Top-1 accuracy of 80.16%.
-In general, GroupNL can accelerate model training because: (i) the non-trainable parameters contained in GroupNL do not need to be transferred between multiple GPUs, reducing the transmission time during model training; (ii) nonlinear transformation operations replace most operations based on sliding window feature map generation, reducing the calculation time during model training.
-GroupNL optimizes the model acceleration in terms of parameter transmission and calculation time.
-Therefore, it can accelerate model training.
+TABLE 16: Module-level Profiling on Raspberry Pi with PyTorch-CPU Backend.
+Module $r$ #Params #FLOPs (M) Avg.Times (ms) FPS ($\uparrow$) Power (W) Energy (Wh)
+Conv - 2.36M 603.98 122.2 8.18 5.003 0.0208
+SineFM Conv 2 1.18M 302.97 97.1 10.3 5.396 0.0183
+4 0.59M 151.81 55.36 18.06 5.285 0.0112
+8 0.30M 76.23 34.86 28.69 5.024 0.0073
+Ghost Conv 2 1.18M 302.58 86.31 11.59 5.262 0.0161
+4 0.59M 151.88 49.95 20.02 5.156 0.0099
+8 0.30M 76.53 31.61 31.63 5.292 0.0072
+GroupNL Conv 2 1.18M 302.06 70.14 14.26 5.005 0.0128
+4 0.59M 151.09 42.32 23.63 4.905 0.0082
+8 0.29M 75.61 28.43 35.17 4.905 0.0063
+Depthwise Conv - 5.12K 1.18 48.71 20.53 6.120 0.0116
+GroupNL Conv (sparse) 2 4.62K 1.25 30.68 32.6 6.096 0.0081
+4 4.63K 1.28 22.65 44.16 5.730 0.0062
+8 4.66K 1.29 18.82 53.14 5.831 0.0056
+Figure 12: Power in Module-level Profiling.
+Figure 13: Energy Consumption in Module-level Profiling.
+TABLE 17: Severity Levels Comparison in EfficientNet-ES Structure (Top: Vanilla; Down: GroupNL) on ImageNet-C.
+Level mCE ($\downarrow$) Corruption Error ($\downarrow$)
+Gauss. Noise Shot Noise Impulse Noise Defocus Blur Glass Blur Motion Blur Zoom Blur Snow Frost Fog Bright. Contrast Elastic Trans. Pixelate JPEG Comp. Speckle Noise Gauss. Blur Spatter Saturate
+Mean 55.033 62.052 63.339 65.818 64.430 73.497 58.716 65.290 59.704 54.999 47.713 30.732 45.922 57.111 60.785 44.755 54.082 60.322 41.914 34.451
+53.958 60.843 61.825 65.581 62.466 72.968 58.113 63.870 57.496 54.063 47.302 30.174 46.991 55.622 58.412 43.345 53.372 58.514 40.574 33.670
+Severity 1 36.260 36.416 36.568 39.068 43.558 46.032 34.822 48.798 40.4 35.408 34.764 27.09 31.504 32.602 36.878 35.164 35.152 33.166 28.818 32.724
+35.599 35.996 36.492 40.402 42.060 44.626 33.856 47.136 38.78 34.846 34.686 26.522 31.214 31.780 35.662 34.420 35.50 32.576 27.910 31.910
+Severity 2 45.169 43.776 46.122 52.036 51.022 59.184 43.698 58.968 60.264 49.080 38.620 28.162 34.250 53.546 43.180 38.784 39.126 46.816 35.982 35.586
+44.314 43.382 45.804 53.748 49.220 58.044 42.622 57.376 57.550 48.358 39.076 27.544 34.332 51.996 41.150 37.810 39.920 45.180 34.616 34.238
+Severity 3 53.486 58.540 60.116 61.688 65.410 83.766 58.614 66.520 56.594 59.804 45.486 29.682 38.798 48.584 54.960 41.356 54.754 61.130 42.196 28.240
+52.404 57.376 58.850 62.140 63.276 83.656 57.852 65.074 54.288 58.764 45.746 29.134 39.670 46.558 52.248 40.040 54.288 58.622 40.414 27.676
+Severity 4 64.426 77.188 81.608 81.218 76.78 87.732 74.308 73.076 66.678 62.122 51.794 32.464 51.208 63.836 77.158 49.130 64.962 73.144 45.842 33.838
+62.845 75.034 78.738 78.738 74.538 87.688 74.090 71.694 63.424 60.934 50.998 31.958 52.784 61.382 72.984 47.538 63.078 70.662 44.416 33.374
+Severity 5 75.826 94.338 92.28 95.080 85.380 90.772 82.136 79.088 74.586 68.580 67.90 36.262 73.848 86.988 91.750 59.340 76.416 87.356 56.734 41.868
+74.628 92.428 89.24 92.878 83.236 90.826 82.146 78.068 73.436 67.412 66.002 35.710 76.956 86.392 90.014 56.918 74.072 85.530 55.514 41.152
+4.4.4 Analysis of Corruption Levels
+We report the results of five corruption levels for EfficientNet-ES in Table 17.
+With very few exceptions (e.g., contrast), GroupNL EfficientNet-ES demonstrates superior robustness across all five severity levels of ImageNet-C corruptions.
+It achieves a lower mCE and outperforms the vanilla model in the vast majority of corruption types at every severity level.
+This consistent improvement highlights GroupNL’s effectiveness in enhancing model robustness against a wide range of real-world image distortions.
 5 Conclusion and Future Work
-This paper presented a robust and low-resource CNN design method, GroupNL.
-In GroupNL, seed feature maps are grouped and nonlinear transformation functions configured with different hyperparameters are used to generate diverse feature maps on demand to reduce the number of parameters and floating-point operations (FLOPs).
-Experimental results show that GroupNL can effectively improve the robustness of the CNN model on multiple tasks while reducing the number of model parameters and FLOPs.
+This paper presented a low-resource and robust CNN design method, GroupNL.
+In GroupNL, seed feature maps are grouped and nonlinear transformation functions configured with different hyperparameters and lightweight tensor manipulation operations are used to generate diverse feature maps on demand to reduce the number of parameters and FLOPs.
+Experimental results show that GroupNL can effectively improve the robustness of CNNs while acceleration the training and inference.
+We will analyze the selection of NLFs from a theoretical perspective in our future work.
+Specifically, we plan to analyze the selection of NLFs from a theoretical perspective based function approximation theory, and through more fine-grained experimental analysis, such as analyzing the distance of the feature maps generated by each NLF.